### v1

# On the Entropy Calibration of Language Models

Abstract
We study the problem of entropy calibration, which asks whether a language model’s entropy over generations matches its log loss on human text. Past work found that models are miscalibrated, with entropy per step increasing (and text quality decreasing) as generations grow longer. This error accumulation is a fundamental problem in autoregressive models, and the standard solution is to truncate the distribution, which improves text quality at the cost of diversity. In this paper, we ask: is miscalibration likely to improve with scale, and is it theoretically possible to calibrate without tradeoffs? To build intuition, we first study a simplified theoretical setting to characterize the scaling behavior of miscalibration with respect to dataset size. We find that the scaling behavior depends on the power law exponent of the data distribution — in particular, for a power law exponent close to 1, the scaling exponent is close to 0, meaning that miscalibration improves very slowly with scale. Next, we measure miscalibration empirically in language models ranging from 0.5B to 70B parameters. We find that the observed scaling behavior is similar to what is predicted by the simplified setting: our fitted scaling exponents for text are close to 0, meaning that larger models accumulate error at a similar rate as smaller ones. This scaling (or, lack thereof) provides one explanation for why we sample from larger models with similar amounts of truncation as smaller models, even though the larger models are of higher quality. However, truncation is not a satisfying solution because it comes at the cost of increased log loss. In theory, is it even possible to reduce entropy while preserving log loss? We prove that it is possible, if we assume access to a black box which can fit models to predict the future entropy of text.
1 Introduction
We study entropy calibration, which asks whether a language model’s entropy over generations matches its log loss on human text. This definition is a natural notion of calibration for generative tasks, and is much more challenging than calibration for classification tasks because the output space is exponentially large. While we will discuss this definition in more detail later, one basic requirement for calibration is that the model’s entropy per step should be stable over the generation.https://github.com/stevenxcao/entropy-calibration
Unfortunately, autoregressive language models are not stable. Entropy calibration was first studied by braverman2020calibration, who found that language models have entropy per step increasing as generations grow longer. This entropy blowup is accompanied by an increase in generation errors, as is also observed in basu2021mirostatneuraltextdecoding. While entropy calibration specifically is not well studied, generation instability more broadly has been the subject of many papers (williams1989learning; Ranzato2015SequenceLT; Welleck2019NeuralTG). From this line of work has emerged a suite of distribution truncation techniques, which have become standard practice in language model sampling (fan-etal-2018-hierarchical; holtzman2020curious; hewitt-etal-2022-truncation). These techniques suppress low probability tokens to improve quality at the cost of diversity (hashimoto-etal-2019-unifying; zhang-etal-2021-trading).
It is not fully satisfying that to make generation stable, we must sacrifice diversity. Diversity is especially important for difficult tasks where we must aggregate multiple answers (wang2024mixtureofagentsenhanceslargelanguage; brown2024largelanguagemonkeysscaling), as well as for synthetic data generation, which has seen a resurgence of interest as the community has begun worrying about running out of internet data (wang-etal-2023-self-instruct; gunasekar2023textbooksneed; maini-etal-2024-rephrasing). Therefore, it is natural to ask: do we expect generation stability to improve with scale? If not, is it at least theoretically possible to calibrate without sacrificing diversity?
To build intuition, we first study a simplified theoretical setting, where instability comes from the fact that the model might generate a token that it saw only a few times during training. This unfamiliar token then derails subsequent steps when it is fed back into the context autoregressively. Drawing on classic results, we calculate a scaling exponent capturing how quickly the probability of generating a rare token decreases with the number of training examples (good-1953-population; karlin-1967-central). We find that this exponent depends on how heavy-tailed the data distribution is: in particular, for power law exponents close to 1, as is typical for human text (zipf-1936-psycho; zipf-1949-human), the scaling exponent is close to 0. Therefore, this setup predicts that stability in generation improves very slowly with scale.
Next, we measure miscalibration empirically in large language models with up to 70B parameters, on three datasets. We find that the observed scaling behavior is similar to what is predicted by the simplified setting: fitting scaling exponents relating calibration error to model size, we find that the exponent for the two text datasets is around 0.05-0.05, meaning that larger models are similarly miscalibrated as smaller ones. On the other hand, for the code dataset, the scaling exponent is around 0.3-0.3, meaning that miscalibration improves moderately with scale. We measure the power law exponent to be around $1$ for the two text datasets, and $1.5$ for the code dataset. Therefore, these findings are consistent with the theory: the code dataset has more quickly decaying tails, so the scaling should indeed be faster. However, further work on more datasets is needed to more strongly establish this relationship between the power law and scaling exponents.
If even large models suffer from error accumulation, why are reasoning and instruction-tuned models able to produce long, coherent outputs? We find that much like distribution truncation, instruction tuning reduces entropy at the cost of increased log loss, with the largest models now having entropy too low. This tradeoff relates to past work which found that alignment degrades model capabilities, a phenomenon known as the alignment tax (ouyang2022training; bai2022helpful; lin-etal-2024-mitigating).
Given that all known mitigations increase the model’s log loss, is it even possible in theory to calibrate without this tradeoff? Drawing on ideas from reinforcement learning theory, we prove that it is possible, if we assume access to a black box which can fit models on the future entropy of text prefixes and attain low test error. Specifically, we describe a polynomial-time calibration procedure that adjusts each candidate token’s probability based on the expected entropy of its continuations. While the resulting procedure is impractical to implement, we prove that it calibrates while preserving log loss, suggesting that generation stability and diversity might be possible to attain simultaneously.
2 Preliminaries
We first review key definitions and properties for entropy calibration, introduced in braverman2020calibration. Our setup is as follows: we are given prompts $X\in\mathcal{X}$ drawn from some prompt distribution $X\sim q$, and responses $Y\in\mathcal{Y}$ drawn from the true conditional distribution $Y\sim p^{*}_{X}$. For example, $X$ might contain a description of a coding task, while $Y$ contains a solution to the task. We then train a language model $\hat{p}:\mathcal{X}\to\Delta^{\mathcal{Y}}$ to fit the true conditional distribution $p^{*}$. We say that $\hat{p}$ is entropy calibrated if its entropy over generations is equal to its log loss:
[…]
where the total/per-step entropy and total/per-step log loss are given by
[…]
To build intuition for this definition, entropy can be thought of as a measure of the model’s uncertainty, which should be calibrated to match the actual loss it incurs on real data. This definition mirrors that of binary calibration, and we derive this connection more formally in Appendix B. Qualitatively, if a model is underconfident, then its generations have too much entropy and appear incoherent; if it is overconfident, then its generations have too little entropy and appear repetitive (braverman2020calibration; basu2021mirostatneuraltextdecoding); see Appendix E for a replication of this finding and Appendix D for examples. Entropy calibration is then the problem of adjusting the entropy to be just right. Empirically, braverman2020calibration found that neural autoregressive language models have entropy too high: entropy per step matches the log loss at earlier steps but increases as the generation grows.
Why does entropy per step grow with the length of the generation? The main problem, as has been observed in empirical work, is that autoregressive language models accumulate error during generation. At training time, models are given input from the true distribution and asked to produce only a single additional token. In contrast, models must generate multiple tokens at deployment time, which they do by producing one token at a time and taking their own production as subsequent input. Therefore, even models with very low single-step error can degrade over multiple steps as they take their own slightly erroneous outputs as input and accumulate errors (see, e.g., Ranzato2015SequenceLT, Welleck2019NeuralTG, holtzman2020curious for error accumulation in language modeling; and daume2009search, ross2010efficient, ross2011reduction for imitation learning). This intuition is formalized in the context of entropy calibration in Corollary 4.2 of braverman2020calibration, which states that for a model with $\varepsilon$ KL divergence to the true distribution, the entropy at step $t$ can deviate as much as $\varepsilon+\sqrt{\varepsilon t}$ from the log loss, growing with $t$.
How does one calibrate the entropy? Unlike binary and multiclass calibration, entropy calibration is challenging because the models have an exponentially large output space. In practice, practitioners use a number of distribution truncation methods, each of which uses a different heuristic to suppress low probability tokens in each generation step. Some standard methods include temperature reduction, top-k sampling (fan-etal-2018-hierarchical), top-p sampling  (holtzman2020curious), and min-p sampling (hewitt-etal-2022-truncation). These methods improve text quality at the cost of diversity (hashimoto-etal-2019-unifying; zhang-etal-2021-trading; pillutla2021mauve; welleck2024decodingmetagenerationinferencetimealgorithms). Following hashimoto-etal-2019-unifying, we define a model’s diversity to be its log loss on reference documents. The intuition behind this definition is that log loss (also known as cross entropy or forward KL) is a coverage metric: to attain low log loss, the model must “cover” as much as the reference distribution as possible. Our goal, then, is to calibrate entropy to match log loss without also causing the log loss to also increase.
In theory, braverman2020calibration show that one can calibrate entropy while preserving log loss via globally normalized temperature scaling, where the adjusted model is given by $\hat{p}_{\tau}(y_{1},...,y_{L})\propto\hat{p}(y_{1},...,y_{L})^{1/\tau}$. Unfortunately, this adjustment is intractable to compute because it involves normalizing over the entire output space. It remains unclear, then, whether this goal is possible in polynomial time. Specifically, we wish to take in a model $\hat{p}$ and produce a calibrated model $\tilde{p}$ with at most $\varepsilon$ entropy calibration error per step, as well as log loss at most that of the original model $\hat{p}$:
[…]
where the entropy calibration error is defined as the difference between the entropy and the log loss:
[…]
3 Intuition: Singleton Mass in Power Law Data
Before putting in the work to develop better calibration algorithms, it is natural to first ask whether we expect miscalibration to automatically improve with scale, as we train larger models on more data. To gain intuition, we first explore this question in a simplified theoretical setting. We define the setup to capture the following hypothesis regarding error accumulation (see, e.g., hewitt-etal-2022-truncation): because the language distribution is heavy-tailed, the model must assign non-zero probability to a large number of rare tokens when fitting the data. However, if it happens to generate one such rare token, the model derails when that token is fed back into the context autoregressively, leading to a jump in entropy. Over many generation steps, then, the model will eventually derail. The degree of instability then depends on the probability of producing a rare token.
Accordingly, our setup is as follows: at training time, the model stores the counts for $m$ tokens drawn i.i.d. from an $\alpha$-power-law distribution $p$ over a vocabulary of size $v$, defined as $p_{i}\propto 1/i^{\alpha}$ for i=1,,vi=1,...,v. The model then generates a sequence token-by-token as follows: if all tokens in context were seen at least twice at training time, the model samples a random token seen during training. But if any token in context was seen only once, the model samples from a high entropy “derailed” distribution instead. This simple stylized setting captures our intuition about error accumulation and lets us study the effect of $\alpha$, representing the heavy-tailedness of the data distribution.
In this setting, the expected entropy per step grows with slope proportional to the probability of emitting a rare token (see Appendix B). Computing the rare token mass in power law data is a classic problem, and we can compute the asymptotic scaling exponent with respect to the number of training examples $m$ as follows (good-1953-population; karlin-1967-central):
Proposition 3.1 (informal).
For $v$ infinite and $m$ large, the per-step probability of generating a singleton, in expectation over draws of the training set, is given by
$\mathbb{E}\frac{K_{m,1}}{m}=C_{\alpha}m^{1/\alpha-1},$
where $C_{\alpha}$ is a constant depending only on $\alpha$, and $K_{m,1}$ is a random variable denoting the number of items seen exactly once in a set of $m$ samples.
We provide the derivation in the appendix. The key takeaway from this proposition is that the derailing probability scales as $m^{1/\alpha-1}$, which is very slow if the power law exponent $\alpha$ is close to $1$, as is typical for text (zipf-1936-psycho; zipf-1949-human). The reason for this slow scaling is that as $m$ increases, there are always more rare items to be sampled from the tail of the distribution. In practice, of course, we are not training unigram models, but the same intuition holds if we posit that semantic concepts in text are similarly heavy tailed: as larger models are trained on more data, there will always be new rare phenomena that they see during training only once. These phenomena are then memorized, and potentially generated at deployment, derailing the model.
While asymptotic analysis gives us a clean expression, we can also estimate the scaling exponent in simulation for finite values of $m$ and $v$ (see Figure 1). We find that the actual slopes are close to the asymptotic expression, up to $m\approx v/3$. We also calculate the power law exponent for our three datasets, finding that it is around 1 for WikiText and WritingPrompts and $1.5$ for CodeContests, which predicts slow scaling for the first two datasets and slow-to-moderate scaling for the third.
Figure 1: Left: the expected total mass of tokens seen exactly once, given m samples from a power law distribution over a vocabulary of size $v$, for three settings of the power law exponent $\alpha=1.0,1.25,1.5$. Their relationship is roughly log-log linear up to $m\approx v/3$, with slope slightly steeper than the asymptotic expression of $1/\alpha-1$. Right: log frequency versus log rank of the top 5000 unigrams in three datasets. The power law exponent $\alpha$, given by the slope of each curve, is close to 11 for WikiText and WritingPrompts, while it is $1.5$ for CodeContests, suggesting that text has heavier tails than code. Together, these plots suggest that the singleton mass should decay more slowly with $m$ for WikiText and WritingPrompts than for CodeContests.
4 Experiments: Miscalibration in Large Language Models
Next, we measure miscalibration empirically in large language models. We study four model families (Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B) (qwen2025qwen25technicalreport), Llama 3 (1B, 3B, 8B, 70B) (grattafiori2024llama3herdmodels), Llama 2 (7B, 13B, 70B) (touvron2023llama2openfoundation), and Pythia (410M, 1.4B, 2.8B, 6.9B, 12B) (biderman2023pythiasuiteanalyzinglarge)) applied to the three datasets listed below. In each setting, we use 5000 examples and limit samples to 1024 tokens; see Appendix C for more experimental details. We primarily study base models because we are interested in the problem of modeling human text; we study the effect of instruction tuning in Section 4.3.
  1. (a)
    WikiText-103 (merity2016pointersentinelmixturemodels): given 128 tokens of context from a Wikipedia passage, the model is tasked with completing the passage.
  2. (b)
    WritingPrompts (fan-etal-2018-hierarchical): given a prompt from r/writingprompts along with 128 tokens of context from a human-written story, the model is tasked with completing the story.
  3. (c)
    CodeContests (li-2022-competition): given a coding problem from one of five websites and 128 tokens of context from a human-written solution, the model is tasked with completing the solution.
4.1 Miscalibration scaling in base models
Figure 2: Log calibration error versus log model size for four model families and three datasets. We find that the scaling laws fit relatively well, suggesting that the relationship between calibration and scale is predictable. Furthermore, while there is variation between model families, the scaling exponents for each dataset are somewhat close to those predicted by theory (WikiText: $0.089$, WritingPrompts: 0.10-0.10, CodeContests: 0.33-0.33), with heavier-tailed datasets having slower scaling.
Past work has found that many model capabilities improve predictably with model size, with task loss and model size following a linear relationship when plotted on a log scale (kaplan2020scalinglawsneurallanguage; hoffmann2022trainingcomputeoptimallargelanguage). We use a similar methodology to study the relationship between entropy miscalibration and model size. If model size and dataset size are scaled proportionally, Proposition 3.1 suggests a scaling law of $\log\text{EntCE}=(1/\alpha-1)\log m+C$, where $\alpha$ is the power law exponent of the data distribution and $m$ is the parameter count. Does the actual data also follow a clean scaling law, and how close is the scaling exponent to that predicted by the simplified setting?
For each model-dataset combination, we compute the model’s calibration error as the difference between its average entropy per generation step and its average log loss on ground truth data. We then plot log calibration error versus log model size, as shown in Figure 2.
First, we find that the linear fit is accurate, suggesting that the relationship between calibration and scale is predictable. Next, we find that the scaling exponents are dataset-dependent: for the older model families (Llama 2 and Pythia), the exponents are around $0.0$ for WikiText and WritingPrompts and 0.2-0.2 for CodeContests, while for the newer model families (Llama 3 and Qwen2.5), the exponents are around 0.13-0.13 for WikiText and WritingPrompts and 0.35-0.35 for CodeContests. Notably, these exponents are somewhat close to what is predicted theoretically (Figure 1): WikiText and WritingPrompts, with power law exponents of $0.918$ and $1.114$, are predicted to have slow scaling exponents of $0.089$ and 0.10-0.10, while CodeContests, with a power law exponent of $1.5$, is predicted to have a moderate scaling exponent of 0.33-0.33. However, future work on more datasets would be needed to more strongly establish the relationship between these exponents empirically.
We speculate that recent model families have better scaling due to changes in their pretraining data mixtures, and especially the addition of a midtraining step with higher quality and less diverse data. However, training details for three out of the four model families (all but Pythia) are not public, and future work with controlled data mixtures would be useful to disentangle the effects of model size, dataset size, and dataset composition.
Overall, these plots suggest that miscalibration in text generation improves very slowly with scale: a scaling exponent of 0.10-0.10 means that to reduce calibration error by a factor of $10$, dataset size must increase by a factor of $10^{10}$.
Figure 3: Entropy for each generation step (solid) and log loss for each token in the ground truth (dashed), for each dataset (columns) and each model family (rows), with models colored by size. Models have entropy much higher than their log loss, with the gap growing with the number of generation steps, a sign of error accumulation. For the text datasets, models of different sizes seem to be similarly miscalibrated, while for code the degree of miscalibration seems to improve with size.
4.2 Entropy over time
To gain a more fine-grained understanding of entropy blowup, we also produce entropy over time plots for each model and each dataset, as shown in Figure 3. Specifically, we plot each model’s entropy at each generation step $t$, averaged over 5000 generated samples. We then compare this curve to the model’s log loss on each token $t$ of a ground truth example, averaged over 5000 examples. Recall from Section 2 that theoretically, for an accurate model, entropy is initially close to log loss, but can deviate as much as $\sqrt{t}$ at the $t$-th step. A calibrated model which does not experience error accumulation should have entropy close to the log loss for all generation steps.
First, we find that for each model and dataset, the log loss is mostly constant or slightly decreasing over time. Past papers use a model’s log loss to estimate the actual entropy of the underlying text, as the former is an upper bound for the latter that grows tighter if the model is more accurate. This part of the plot replicates past findings that the entropy per step of human text is stable over time, also known as the entropy rate constancy principle (genzel-charniak-2002-entropy; verma-etal-2023-revisiting).
On the other hand, unlike human text, the entropy per step of model generations is not stable and instead increases over time. The lack of scaling shown quantitatively in the previous subsection is reflected visually in Figure 3, with larger models having entropy growing at similar rates as smaller models for WikiText and WritingPrompts (the left and middle columns). For CodeContests, the slopes decrease with model size, visually confirming that there is slow-to-moderate scaling.
Figure 4: Entropy calibration error versus log loss for base Qwen2.5 (1.5B, 7B, 72B) compared to the instruction-tuned versions, along with various temperature settings (please see Appendix E for all model sizes). Positive calibration error means that the model’s entropy is higher than its log loss, while negative means that its entropy is lower than its log loss. We find that each modification of the base model reduces entropy while increasing log loss, calibrating at the cost of diversity.
4.3 Calibration-diversity tradeoffs
In this subsection, we seek to better understand how distribution truncation and post-training affect entropy. For each model-dataset combination (excluding Pythia, which has no instruction-tuned version), we compare the model with temperature $1.0$ to that with temperature $0.95$, $0.9$, $0.85$, or $0.8$, as well as to the instruction-tuned version of the model. We then plot entropy calibration error against log loss, where each setting of the model is one point on the plot, as shown in Figure 4.
First, we find that reducing temperature below $1$ reduces entropy but increases log loss, replicating similar findings in past work (hashimoto-etal-2019-unifying). Furthermore, the temperature attaining zero calibration error is similar across model sizes, which makes sense given that they are similarly miscalibrated. We find that instruction tuning also reduces entropy while increasing log loss, which is consistent with past work showing that instruction tuning harms diversity (ghosh2024closer). Unlike temperature scaling, the magnitude of the effect varies across model sizes, with larger models experiencing both a larger reduction in entropy and larger increase in log loss. However, this pattern is not robust across model families (see Appendix E). Further work with more controlled instruction tuning would be necessary to explore this relationship further. These experiments reconcile our previous finding, that even large models accumulate errors, with the fact that in practice, one can use truncation or post-training to generate long, coherent pieces of text. The tradeoff is that each of these mitigations comes at the cost of diversity.
5 Theory: Future Entropy Scaling
If all known mitigations increase log loss, is it even possible in theory to calibrate without this tradeoff? In this section, we provide evidence that this tradeoff is not inevitable: given the assumption that there exists a procedure to fit regression models that generalize to i.i.d. test data, we show that there exists a tractable, albeit impractical, procedure that calibrates while preserving log loss.
Algorithm 1 Future entropy scaling
Inputs: model $\hat{p}$, length $T$, vocab $\mathcal{V}$, future entropy fitting algorithm $\mathcal{A}$, future entropy dataset size $n$, sample size $m$, prompt distribution $q$, true conditional distribution $p^{*}$, optimization tolerance $\varepsilon$
  1. 1:
    Initialize $\alpha_{1}=...=\alpha_{T}=0$, $\hat{f}_{2}=...=\hat{f}_{T+1}=0$.
  2. 2:
    For t=T,,1t=T,...,1:
  3. 3:
      Choose $\alpha_{t}$ to minimize expected log loss at step $t$, until the gradient is at most $\varepsilon$:
    $\alpha_{t}=\operatorname*{argmin}_{\alpha_{t}^{\prime}}\mathcal{L}_{t}\left(p^{*}\ \|\ \hat{p}_{\alpha^{\prime},\hat{f}}^{\text{(ent)}}\right)$
      where $\alpha^{\prime}=(0,...,0,\alpha_{t}^{\prime},\alpha_{t+1},...,\alpha_{T})$. ($\mathcal{L}_{t}$: Equation 3, $\hat{p}_{\alpha^{\prime},\hat{f}}^{\text{(ent)}}$: Equation 9).
  4. 4:
      Fit the future entropy predictor $\hat{f}_{t}$ as follows:
  5. 5:
       Sample prefixes $\left(X^{(i)},Y_{<t-1}^{(i)}\right)_{i=1}^{n}\sim(q,p^{*})$.
  6. 6:
       For each token $v\in\mathcal{V}$, compute labels $(h^{(i,v)})_{i=1}^{n}$ by passing each prefix $\left(X^{(i)},\left[Y_{<t-1}^{(i)},v\right]\right)$ into Algorithm 2, along with inputs $\hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}$, $T$, $m$.
  7. 7:
       Fit one future entropy predictor for each token $v$, setting $\hat{f}_{t}(X,[Y_{<t-1},v])=\hat{f}_{t,v}(X,Y_{<t-1})$, where each $\hat{f}_{t,v}$ is the output $\mathcal{A}\left\{\left(X^{(i)},Y_{<t-1}^{(i)},h^{(i,v)}\right)_{i=1}^{n}\right\}$.
  8. 8:
    Return $(\alpha_{1},...,\alpha_{T}),(\hat{f}_{2},...,\hat{f}_{T+1})$.
Algorithm 2 Future entropy estimation via sampling
Inputs: model $\hat{p}$, length $T$, prefix $(X,Y_{\leq t})$, samples $m$
  1. 1:
    Sample $m$ trajectories from the model: $\left(\hat{Y}_{t+1}^{(i)},...,\hat{Y}_{T}^{(i)}\right)_{i=1}^{m}\overset{\textrm{i.i.d.}}{\sim}\hat{p}_{X}(\hat{Y}_{>t}\mid Y_{\leq t})$.
  2. 2:
    Return $\hat{H}=\frac{1}{m}\sum_{i=1}^{m}\sum_{s=t+1}^{T}-\log\hat{p}_{X}(\hat{Y}_{s}^{(i)}\mid\hat{Y}_{<s}^{(i)})$.
5.1 Definitions
For a model $\hat{p}_{X}(Y_{1},...,Y_{T})$ mapping a prompt $X$ to a distribution $\Delta^{\mathcal{Y}}$ over the output space $\mathcal{Y}=\mathcal{V}^{T}$, let the future entropy of the prefix $(X,Y_{\leq t})$ be given by
$H_{\hat{p}_{X}}(Y_{>t}\mid Y_{\leq t})=\sum_{Y_{>t}}-\hat{p}_{X}(Y_{>t}\mid Y_{\leq t})\log\hat{p}_{X}(Y_{>t}\mid Y_{\leq t}).$ (7)
Given a prefix $Y_{\leq t}$, this expression computes the model’s entropy over the remaining generation $Y_{>t}$. We can then define the future entropy adjusted model, for parameters $\alpha=(\alpha_{1},...,\alpha_{T})\in\mathbb{R}^{T}$, as
$\hat{p}_{\alpha;X}^{\text{(ent)}}(Y_{t}\mid Y_{<t})$ $\propto\exp\left\{(1+\alpha_{t})\log\hat{p}_{X}(Y_{t}\mid Y_{<t})-\alpha_{t}H_{\hat{p}_{\alpha;X}^{\text{(ent)}}}(Y_{>t}\mid Y_{\leq t})\right\}.$ (8)
This expression adjusts each candidate token’s probability based on what the future entropy would be if that token were chosen. The calibration procedure then involves fitting models to predict the future entropy of prefixes, and choosing the weights $\alpha_{t}$ to calibrate the model (Algorithm 1).
5.2 Assumptions
For a distribution $\hat{p}$ that can be tractably sampled from, we can take a Monte Carlo estimate to compute the future entropy, which concentrates because entropy is bounded (Algorithm 2). However, we cannot assume that $\hat{p}_{\alpha;X}^{\text{(ent)}}$ can be tractably sampled from, so we cannot compute its future entropy naively. Instead, we will use our assumed model fitting procedure to iteratively replace each intractable future entropy term $H_{\hat{p}_{\alpha;X}^{\text{(ent)}}}(Y_{>t}\mid Y_{\leq t})$ with a tractable fitted model $\hat{f}(X,Y_{\leq t})$, leading to the following approximate future entropy adjustment:
$\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{t}\mid Y_{<t})$ $\propto\exp\left\{(1+\alpha_{t})\log\hat{p}_{X}(Y_{t}\mid Y_{<t})-\alpha_{t}\hat{f}_{t+1}(X,Y_{\leq t})\right\}.$ (9)
Then, we can initialize $\alpha=(0,...,0)$ and first fit $\alpha_{T}$ for the last generation step. Next, now that the last generation step is calibrated, we can fit future entropy model $\hat{f}_{T}$, taking in length T1T-1 prefixes and predicting the entropy at step $T$. Given $\hat{f}_{T}$, we can then fit $\alpha_{T-1}$, calibrating the second to last step. This procedure proceeds from t=T,,1t=T,...,1, resulting in a calibrated model.
The future entropy model fitting relies on the following assumption, which states that we can fit regression models that attain good i.i.d. test error:
Assumption 5.1.
Let $\left(X^{(i)},Y_{\leq t}^{(i)},h^{(i)}\right)_{i=1}^{n}$ be a dataset with inputs $X^{(i)}\overset{\textrm{i.i.d.}}{\sim}q$ and $Y_{\leq t}^{(i)}\sim p^{*}_{X^{(i)}}$, along with noisy labels $h^{(i)}$. Furthermore, suppose each noisy label is given by $h^{(i)}=f^{*}(X^{(i)},Y_{\leq t}^{(i)})+\varepsilon_{i}$ for the true label $f^{*}(X^{(i)},Y_{\leq t}^{(i)})\in\mathbb{R}$ and noise $\varepsilon_{i}\overset{\textrm{i.i.d.}}{\sim}\mathcal{E}$, where $\mathcal{E}$ is a mean-zero noise distribution bounded by $\sigma$. Then, there exists a polynomial time algorithm $\mathcal{A}$ which takes in a dataset of size $n\in\text{poly}(\sigma,\delta)$ and outputs a fitted model $\hat{f}$ attaining at most $\delta$ test error:
$\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}|\hat{f}(X,Y_{\leq t})-f^{*}(X,Y_{\leq t})|\leq\delta.$
Empirically, neural networks can fit almost anything while attaining low test error in distribution, and the future entropy prediction problem described here is not particularly complex, involving mapping a set of tokens to a single bounded scalar. However, future empirical work is needed to determine how accurately a large neural model can predict the future entropy.
5.3 Main theorem
With this assumption, we now state the main result:
Theorem 5.2.
[…]
$\left|\text{EntCE}\left(p^{*}\ \|\ \hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}\right)\right|$ $\leq 2T\delta+\sum_{t=1}^{T}(1+\alpha_{t})\varepsilon,$
[…]
This theorem tells us that if each future entropy predictor has error $\delta$ and we choose each $\alpha_{t}$ to be an $\varepsilon$-stationary point with respect to the log loss, the calibrated model will have entropy within $O(\delta+\varepsilon)$ of its log loss at each time step, and its log loss will be better than that of the original model.
Why does future entropy preserve log loss? Future entropy adjustment can be derived as a first-order approximation of globally normalized temperature adjustment; we provide this derivation in Appendix B, along with a derivation in the MaxEnt RL framework (ziebart-2008-maximum). Global temperature adjustment attains calibration as long as the gradient of the log loss with respect to temperature is small (braverman2020calibration), which is a first-order condition. Then, intuitively, a first-order approximation of global temperature scaling should preserve this property.
The procedure described in Algorithm 1 is not practical to implement, as one would need to a fit a separate future entropy predictor for each generation step and each candidate token, each of which involves a slow data collection process based on a repeated sampling. Nonetheless, the existence of such an algorithm provides evidence that log loss tradeoffs are not inevitable in entropy calibration, despite the output space being exponentially large. One other point to note is that our analysis holds for any approximation of the future entropy that attains error $\delta$, with worse approximations just weakening the calibration error guarantee. For example, one could use the one-step future entropy (braverman2020calibration), or truncate to $k$ steps instead. We hope that our theory, which establishes future entropy as the target to approximate, guides future work to achieve better quality-diversity tradeoffs than existing approaches.
6 Additional Related Work
Calibration is most commonly studied in binary and multiclass classification, with some classic algorithms including binning, Platt scaling, and isotonic regression (PlattProbabilisticOutputs1999; zadrozny2002transforming; guo2017on; kumar2019verified). In the language modeling setting, liang2023holisticevaluationlanguagemodels evaluate the calibration of language models prompted to perform a wide range of classification tasks, finding that models are almost always miscalibrated and overconfident. In such a setting, one can simply apply standard calibration techniques to adjust the model’s outputted probabilities. More challenging is linguistic calibration, where models appear overconfident in the language they use to answer a question. To address this problem, past works propose techniques based on controllable generation and reinforcement learning (mielke-etal-2022-reducing; band2024linguistic). Finally, the term “calibration” is also used to describe the procedure of eliminating the model’s innate bias toward certain tokens when doing in-context learning, to improve task performance (pmlr-v139-zhao21c). All of these settings are distinct from our setting, which studies the calibration of a model’s entropy over an entire generation, and whose related work we discuss in Section 2.
7 Conclusion
We find both theoretically and experimentally that entropy miscalibration improves very slowly with scale. Furthermore, while all current methods calibrate at the cost of diversity, we provide theoretical evidence that this tradeoff can be avoided. Therefore, given recent community interest in test-time scaling and synthetic data, both for which diversity is centrally important, we are excited about work which seeks to attain both generation stability and diversity simultaneously.
Acknowledgements
We would like to thank Rishi Bommasani, Sarah Cen, Irena Gao, Konwoo Kim, Suhas Kotha, John Thickstun, and anonymous reviewers for useful conversations about the paper. GV is currently affiliated with OpenAI but did this work while at Stanford. GV and SC were supported by NSF Award AF-2341890 and the Simons Foundation Investigator Award, PL and SC were supported by the Open Philanthropy Project Award, and SC was supported by the NSF Graduate Research Fellowship Program.
Appendix A Proofs
Recall notation: we are given prompts $X\in\mathcal{X}$ drawn from some prompt distribution $X\sim q$, and responses $Y\in\mathcal{Y}$ drawn from the true conditional distribution $Y\sim p^{*}_{X}$ for $p^{*}_{X}\in\Delta^{\mathcal{Y}}$. For simplicity, let $\mathcal{Y}$ be the set $\mathcal{V}^{T}$ of length $T$ strings over a vocabulary $\mathcal{V}$. We then train a language model $\hat{p}:\mathcal{X}\to\Delta^{\mathcal{Y}}$ to fit the true conditional distribution $p^{*}$.
Algorithm 3 Future entropy scaling
Inputs: model $\hat{p}$, length $T$, vocab $\mathcal{V}$, future entropy fitting algorithm $\mathcal{A}$, future entropy dataset size $n$, sample size $m$, prompt distribution $q$, true conditional distribution $p^{*}$, optimization tolerate $\varepsilon$
  1. 1:
    Initialize $\alpha_{1}=...=\alpha_{T}=0$, $\hat{f}_{2}=...=\hat{f}_{T+1}=0$.
  2. 2:
    For t=T,,1t=T,...,1:
  3. 3:
      Choose $\alpha_{t}$ to minimize expected log loss at step $t$, until the gradient is at most $\varepsilon$:
    $\alpha_{t}=\operatorname*{argmin}_{\alpha_{t}^{\prime}}\mathcal{L}_{t}\left(p^{*}\ \|\ \hat{p}_{\alpha^{\prime},\hat{f}}^{\text{(ent)}}\right)$
      where $\alpha^{\prime}=(0,...,0,\alpha_{t}^{\prime},\alpha_{t+1},...,\alpha_{T})$.
      ($\mathcal{L}_{t}$: Equation 3, $\hat{p}_{\alpha^{\prime},\hat{f}}^{\text{(ent)}}$: Equation 9)
  4. 4:
      Fit the future entropy predictor $\hat{f}_{t}$ as follows:
  5. 5:
       Sample prefixes $\left(X^{(i)},Y_{<t-1}^{(i)}\right)_{i=1}^{n}$ with $X^{(i)}\overset{\textrm{i.i.d.}}{\sim}q,\ Y_{<t-1}^{(i)}\sim p^{*}_{X^{(i)}}$.
  6. 6:
       For each token $v\in\mathcal{V}$, compute labels $(h^{(i,v)})_{i=1}^{n}$ by passing each prefix $\left(X^{(i)},\left[Y_{<t-1}^{(i)},v\right]\right)$ into Algorithm 2, along with inputs $\hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}$, $T$, $m$.
  7. 7:
       Fit one future entropy predictor for each token $v$, setting $\hat{f}_{t}(X,[Y_{<t-1},v])=\hat{f}_{t,v}(X,Y_{<t-1})$, where each $\hat{f}_{t,v}$ is the output $\mathcal{A}\left\{\left(X^{(i)},Y_{<t-1}^{(i)},h^{(i,v)}\right)_{i=1}^{n}\right\}$.
  8. 8:
    Return $(\alpha_{1},...,\alpha_{T}),(\hat{f}_{2},...,\hat{f}_{T+1})$.
We say that $\hat{p}$ is entropy calibrated if its entropy over generations is equal to its log loss, in expectation over the prompt:
[…]
where the total entropy and total log loss are given by
[…]
We can also write the per-step entropy and per-step log loss as
[…]
Let the total entropy calibration error be given by
[…]
$=\left|\sum_{t=1}^{T}H_{t}(\hat{p})-\mathcal{L}_{t}(\hat{p}\ \|\ p^{*})\right|.$ (15)
Our goal will be to calibrate the model $\hat{p}$ while preserving its log loss, which we will do by the following adjustment:
$\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{t}\mid Y_{<t})$ $\propto\exp\bigg\{(1+\alpha_{t})\log\hat{p}_{X}(Y_{t}\mid Y_{<t})-\alpha_{t}\hat{f}_{t+1}(X,Y_{\leq t})\bigg\},$ (16)
where $\alpha_{1},...,\alpha_{t}$ denote the adjustment parameters, and $\hat{f}_{2},...,\hat{f}_{T+1}$ denote future entropy predictors (with $\hat{f}_{T+1}=0$), whose goal is to approximate the future entropy. Using Algorithm 3 (copied from Algorithm 1 for convenience) to fit each $\alpha_{t},\hat{f}_{t}$, we show the following result:
Theorem A.1.
[…]
$\left|\text{EntCE}\left(p^{*}\ \|\ \hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}\right)\right|$ $\leq 2T\delta+\sum_{t=1}^{T}(1+\alpha_{t})\varepsilon,$
[…]
The proof proceeds as follows: first, recall that for each step $t$, we choose $\alpha_{t}$ to minimize $\mathcal{L}_{t}\left(p^{*}\ \|\ \hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}\right)$. The first lemma will show that if the future entropy predictor $\hat{f}_{t+1}$ fitted in the previous iteration has at most $\delta$ error (in expectation over $Y_{<t}$ and uniformly over $Y_{t}\in\mathcal{V}$), then this choice of $\alpha_{t}$ produces a calibration-like guarantee.
Lemma A.2.
Suppose that $\alpha_{t}$ is an $\varepsilon$-stationary point with respect to $\mathcal{L}_{t}$:
$\left|\frac{d}{d\alpha_{t}}\mathcal{L}_{t}\left(p^{*}\ \|\ \hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}\right)\right|\leq\varepsilon,$
and that the future entropy predictor $\hat{f}_{t+1}$ attains at most $\delta$ error, in expectation over $Y_{<t}$ and uniformly over $Y_{t}\in\mathcal{V}$:
$\max_{Y_{t}\in V}\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\left|\hat{f}_{t+1}(X,Y_{\leq t})-H_{\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}}(Y_{>t}\mid Y_{\leq t})\right|\leq\delta.$
Then, we have the following calibration guarantee:
$\Bigg|$ $\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{>t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{\leq t})}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{\leq t},\hat{Y}_{>t})\right]$
- $\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{\geq t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{<t})}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{<t},\hat{Y}_{\geq t})\right]\Bigg|\leq(1+\alpha_{t})\varepsilon+2\delta.$
This bound can be thought of as a partial calibration guarantee in the sense that it allows us to swap $Y_{t}\sim p^{*}$ and $\hat{Y}_{t}\sim\hat{p}^{\text{(ent)}}_{\alpha,\hat{f}}$ in the expectation.
To show that Algorithm 3 improves log loss, note that each $\alpha_{t}$ is initialized to 0, so the initial model $\hat{p}^{\text{(ent)}}_{\alpha,\hat{f}}$ is equal to $\hat{p}$. Then, it suffices to show that each iteration of the algorithm improves the log loss, relative to the previous iteration. This statement is true by the following lemma, which states that at each step $t$ in the algorithm, optimizing $\mathcal{L}_{t}$ is equivalent to optimizing the overall log loss $\mathcal{L}$:
Lemma A.3.
Let $\alpha_{t+1},...,\alpha_{T}$ be set arbitrarily, and let $\alpha_{1}=...=\alpha_{t-1}=0$. Also, let $\hat{f}$ be set arbitrarily. Then,
$\operatorname*{argmin}_{\alpha_{t}^{\prime}}\mathcal{L}_{t}\left(p^{*}\ \|\ \hat{p}_{\alpha^{\prime},\hat{f}}^{\text{(ent)}}\right)=\operatorname*{argmin}_{\alpha_{t}^{\prime}}\mathcal{L}\left(p^{*}\ \|\ \hat{p}_{\alpha^{\prime},\hat{f}}^{\text{(ent)}}\right),$
where $\alpha^{\prime}=(0,...,0,\alpha_{t}^{\prime},\alpha_{t+1},...,\alpha_{T})$.
The final lemma involves showing that each future entropy predictor outputted by the algorithm attains low error and satisfies the condition in Lemma A.2. This lemma relies on the fact that the future entropy $H_{\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}}(Y_{>t}\mid Y_{\leq t})$ only depends on $\alpha_{t+1},...,\alpha_{T}$ and $\hat{f}_{t+2},...,\hat{f}_{T+1}$, because it only involves generation steps $t+1$ and onward. Therefore, after $\alpha_{t+1}$ is chosen, the generation process is fixed for steps $t+1$ and onward, so we can fit a future entropy predictor over those steps despite not having yet chosen $\alpha_{1},...,\alpha_{t}$. These facts, along with the black box fitting procedure provided in Assumption 5.1, lead to the following lemma:
Lemma A.4.
For any $\alpha=(\alpha_{1},...,\alpha_{T})$ and $\hat{f}=(\hat{f}_{2},...,\hat{f}_{T+1})$, and for some fixed $t$, let $\alpha^{\prime}=(0,...,0,\alpha_{t},...,\alpha_{T})$ and $\hat{f}^{\prime}=(0,...,0,\hat{f}_{t+1},...,\hat{f}_{T+1})$ be the results of zeroing out the first t1t-1 entries of $\alpha$ and $\hat{f}$. Then, we have that
$H_{\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}}(Y_{>t-1}\mid Y_{\leq t-1})=H_{\hat{p}_{\alpha^{\prime},\hat{f}^{\prime};X}^{\text{(ent)}}}(Y_{>t-1}\mid Y_{\leq t-1})$
for all $Y_{\leq t-1}$. Furthermore, suppose that Assumption 5.1 holds, and let $\mathcal{D}=\left(X^{(i)},Y_{<t-1}^{(i)},h^{(i,v)}\right)_{i=1}^{n}$ be a dataset with
$h^{(i,v)}$ $=H_{\hat{p}_{\alpha^{\prime},\hat{f}^{\prime};X}^{\text{(ent)}}}\left(Y_{>t-1}\mid\left[Y_{<t-1}^{(i)},v\right]\right)+\varepsilon_{i,v}$
$X^{(i)}$ $\overset{\textrm{i.i.d.}}{\sim}q,\ Y_{<t-1}^{(i)}\sim p^{*}_{X^{(i)}},\varepsilon_{i,v}\sim\mathcal{E}$
for some token $v\in\mathcal{V}$, dataset size $n=\text{poly}(T\log\mathcal{V},\delta)$, and some mean-zero noise distribution $\mathcal{E}$ bounded by $T\log\mathcal{V}$. Then, letting $\mathcal{A}$ denote the black box fitting procedure in Assumption 5.1, we have that $\hat{f}_{t,v}=\mathcal{A}(\mathcal{D})$ satisfies
$\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t-1}\sim p^{*}_{X}}\left|\hat{f}_{t,v}(X,Y_{<t-1})-H_{\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}}(Y_{>t-1}\mid[Y_{<t-1},v])\right|\leq\delta.$
We use these lemmas to prove Theorem A.1 as follows:
Proof of Theorem A.1.
Let $\alpha=(\alpha_{1},...,\alpha_{T})$ and $\hat{f}=(\hat{f}_{2},...,\hat{f}_{T+1})$ denote the outputs of the algorithm. It suffices to show the following three inequalities for all $t$:
  1. (a)
    Prediction error bound: the predictor $\hat{f}_{t+1}$ satisfies
    $\max_{Y_{t}\in V}\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\left|\hat{f}_{t+1}(X,Y_{\leq t})-H_{\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}}(Y_{>t}\mid Y_{\leq t})\right|\leq\delta.$
  2. (b)
    Calibration bound: after iteration $t$ of the algorithm, we have
    $\Bigg|$ $\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{>t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{\leq t})}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{\leq t},\hat{Y}_{>t})\right]$
    - $\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{\geq t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{<t})}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{<t},\hat{Y}_{\geq t})\right]\Bigg|\leq(1+\alpha_{t})\varepsilon+2\delta.$
  3. (c)
    Log loss improvement: letting $\alpha^{(t)}=(0,...,0,\alpha_{t},...,\alpha_{T})$ and $\hat{f}^{(t)}=(0,...,0,\hat{f}_{t+1},...,\hat{f}_{T+1})$, we have
    $\mathcal{L}\left(p^{*}\ \|\ \hat{p}_{\alpha^{(t)},\hat{f}^{(t)}}^{\text{(ent)}}\right)\leq\mathcal{L}\left(p^{*}\ \|\ \hat{p}_{\alpha^{(t+1)},\hat{f}^{(t+1)}}^{\text{(ent)}}\right).$
The theorem follows from combining these inequalities for all $t$: first, to show that log loss improves, it suffices to apply inequality (c) (log loss improvement) for all $t$, where $\hat{p}_{\alpha^{(1)},\hat{f}^{(1)}}^{\text{(ent)}}=\hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}$ and $\hat{p}_{\alpha^{(T+1)},\hat{f}^{(T+1)}}^{\text{(ent)}}=\hat{p}$. Similarly, the calibration result follows from applying inequality (b) (calibration bound) for all $t$ with triangle inequality:
$\left|\text{EntCE}\left(p^{*}\ \|\ \hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}\right)\right|$ $=\left|\mathbb{E}_{X\sim q}\mathbb{E}_{Y\sim p^{*}_{X}}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y)\right]-\mathbb{E}_{X\sim q}\mathbb{E}_{Y\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y)}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y)\right]\right|$
$=\Bigg|\sum_{t=1}^{T}\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{>t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{\leq t})}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{\leq t},\hat{Y}_{>t})\right]$
$\qquad-\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{\geq t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{<t})}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{<t},\hat{Y}_{\geq t})\right]\Bigg|$
$\leq\sum_{t=1}^{T}[(1+\alpha_{t})\varepsilon+2\delta].$
Then, showing inequalities (a), (b), and (c) for all $t$ completes the proof. First, note that if inequality (a) (prediction error bound) holds for all $t$, then the other two inequalities follow directly from the lemmas: inequality (a) ensures the condition in Lemma A.2 is satisfied, directly proving inequality (b) (calibration bound). Inequality (c) (log loss improvement) follows from the fact that $\alpha_{T}$ is chosen via $\operatorname*{argmin}_{\alpha_{T}^{\prime}}\mathcal{L}_{T}\left(p^{*}\ \|\ \hat{p}_{\alpha^{\prime},\hat{f}}^{\text{(ent)}}\right)$, which by Lemma A.3 is equivalent to minimizing the overall log loss $\mathcal{L}\left(p^{*}\ \|\ \hat{p}_{\alpha^{\prime},\hat{f}}^{\text{(ent)}}\right)$.
To show inequality (a) (prediction error bound), first note that for $t=T$, it holds trivially because the future entropy is 0. For t=1,,T1t=1,...,T-1, the prediction error bound follows directly from applying Lemma A.4 for each $\hat{f}_{t+1,v}$ for $v\in\mathcal{V}$, where each noisy future entropy label computed via parallel sampling (Algorithm 2) has mean equal to the future entropy and is bounded by $(T-t)\log\mathcal{V}$. ∎
The proofs of the three lemmas proceed as follows:
Proof of Lemma A.2.
Taking the derivative of the log loss $\mathcal{L}_{t}$ with respect to $\alpha_{t}$, we have
$\varepsilon$ $\geq\left|\frac{d}{d\alpha_{t}}\mathcal{L}_{t}\left(p^{*}\ \|\ \hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}\right)\right|$
$=\left|\frac{d}{d\alpha_{t}}\mathbb{E}_{X\sim q}\mathbb{E}_{Y\sim p^{*}_{X}}[-\mathbb{1}_{Y_{t}}(\cdot)^{T}\log\text{softmax}((1+\alpha_{t})\log\hat{p}_{X}(\cdot\mid Y_{<t})-\alpha_{t}\hat{f}_{t+1}(X,[Y_{<t},\cdot]))]\right|$
$=\left|\mathbb{E}_{X\sim q}\mathbb{E}_{Y\sim p^{*}_{X}}\left[-\left(\mathbb{1}_{Y_{t}}(\cdot)-\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{<t})\right)^{T}(\log\hat{p}_{X}(\cdot\mid Y_{<t})-\hat{f}_{t+1}(X,[Y_{<t},\cdot]))\right]\right|,$
where we use $f(\cdot)\in\mathbb{R}^{|\mathcal{V}|}$ to denote the vector $f(v)]_{v\in\mathcal{V}}$, the indicator function is given by $\mathbb{1}_{Y_{t}}(v)=1$ iff $Y_{t}=v$, and $\text{softmax}:\mathbb{R}^{|\mathcal{V}|}\to\mathbb{R}^{|\mathcal{V}|}$ applies the softmax operation, which exponentiates each entry and then normalizes the vector by its sum. Splitting this term into two expectations results in the expression
$=\Bigg|\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}\left[-(\log\hat{p}_{X}(Y_{t}\mid Y_{<t})-\hat{f}_{t+1}(X,Y_{\leq t}))\right]$
$-\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{<t})}\left[-(\log\hat{p}_{X}(\hat{Y}_{t}\mid Y_{<t})-\hat{f}_{t+1}(X,[Y_{<t},\hat{Y}_{t}]))\right]\Bigg|,$
where the two terms differ in whether $Y_{t}\sim p^{*}$ or $\hat{Y}_{t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}$. Multiplying both sides by $(1+\alpha_{t})$, we have
$(1+\alpha_{t})\varepsilon$ $\geq\Bigg|\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}\left[-(1+\alpha_{t})(\log\hat{p}_{X}(Y_{t}\mid Y_{<t})-\hat{f}_{t+1}(X,Y_{\leq t}))\right]$
$-\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{<t})}\left[-(1+\alpha_{t})(\log\hat{p}_{X}(\hat{Y}_{t}\mid Y_{<t})-\hat{f}_{t+1}(X,[Y_{<t},\hat{Y}_{t}]))\right]\Bigg|.$
Next, noticing that both expressions include unnormalized logits for the distribution $p_{\alpha,\hat{f};X}^{\text{(ent)}}$ applied to either $Y_{t}$ or $\hat{Y}_{t}$, we can subtract the same normalizing constant from both expressions, resulting in
$=\Bigg|\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}\left[-\left(\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{t}\mid Y_{<t})-\hat{f}_{t+1}(X,Y_{\leq t})\right)\right]$
$-\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{<t})}\left[-\left(\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\hat{Y}_{t}\mid Y_{<t})-\hat{f}_{t+1}(X,[Y_{<t},\hat{Y}_{t}])\right)\right]\Bigg|.$
Next, to turn each conditional probability into a joint probability, we can add $\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{<t})\right]$ to both expressions:
$=\Bigg|\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}\left[-\left(\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{t},Y_{<t})-\hat{f}_{t+1}(X,Y_{\leq t})\right)\right]$
$-\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{<t})}\left[-\left(\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\hat{Y}_{t},Y_{<t})-\hat{f}_{t+1}(X,[Y_{<t},\hat{Y}_{t}])\right)\right]\Bigg|.$
At this point, we can use the fact that $\hat{f}_{t+1}$ is within $\delta$ of the future entropy (in expectation over $X\sim q,\ Y_{<t}\sim p^{*}_{X}$ and uniformly over $Y_{t}$) to produce the bound
$(1+\alpha_{t})\varepsilon$ $+2\delta\geq\Bigg|\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}\left[-\left(\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(Y_{t},Y_{<t})-H_{\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}}(Y_{>t}\mid Y_{\leq t})\right)\right]$
$-\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{<t})}\left[-\left(\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\hat{Y}_{t},Y_{<t})-H_{\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}}(Y_{>t}\mid[Y_{<t},\hat{Y}_{t}])\right)\right]\Bigg|.$
Finally, note that the future entropy is defined as
$H_{\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}}(Y_{>t}\mid Y_{\leq t})$ $=\mathbb{E}_{\hat{Y}_{>t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{\leq t})}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\hat{Y}_{>t}\mid Y_{\leq t})\right],$
which we can substitute into the previous equation to produce the desired result:
$(1+\alpha_{t})\varepsilon+2\delta$ $\geq\Bigg|\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{\leq t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{>t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{\leq t})}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\hat{Y}_{>t},Y_{t},Y_{<t})\right]$
$-\mathbb{E}_{X\sim q}\mathbb{E}_{Y_{<t}\sim p^{*}_{X}}\mathbb{E}_{\hat{Y}_{\geq t}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{<t})}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\hat{Y}_{>t},\hat{Y}_{t},Y_{<t})\right]\Bigg|.$
Proof of Lemma A.3.
Let $t$ denote the time step of interest. Writing the full log loss as a sum over $s$, we have
$\mathcal{L}\left(p^{*}\ \|\ \hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}\right)$ $=\sum_{s=1}^{T}\mathcal{L}_{s}\left(p^{*}\ \|\ \hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}\right).$
By the definition of $\hat{p}_{\alpha,\hat{f}}^{\text{(ent)}}$, the $t$-th parameter $\alpha_{t}$ has no effect on summands $s\neq t$. Therefore, optimizing the entire sum is equivalent to optimizing only the summand corresponding to $s=t$, proving the desired result. ∎
Proof of Lemma A.4.
First, to show that
$H_{\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}}(Y_{>t-1}\mid Y_{\leq t-1})=H_{\hat{p}_{\alpha^{\prime},\hat{f}^{\prime};X}^{\text{(ent)}}}(Y_{>t-1}\mid Y_{\leq t-1})$
where $\alpha^{\prime},\hat{f}^{\prime}$ are the results of zeroing out the first t1t-1 entries of $\alpha,\hat{f}$, we can simply write out the definition of the future entropy:
$H_{\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}}(Y_{>t-1}\mid Y_{\leq t-1})$ $=\mathbb{E}_{\hat{Y}_{>t-1}\sim\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\cdot\mid Y_{\leq t-1})}\left[-\log\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\hat{Y}_{>t-1}\mid Y_{\leq t-1})\right],$
where we can write out the probability as
$\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\hat{Y}_{>t-1}\mid Y_{\leq t-1})$ $=\prod_{s=t}^{T}\hat{p}_{\alpha,\hat{f};X}^{\text{(ent)}}(\hat{Y}_{s}\mid Y_{\leq t-1},\hat{Y}_{t,...,s-1})$
$=\prod_{s=t}^{T}\mathbb{1}_{\hat{Y}_{s}}^{T}\text{softmax}\bigg((1+\alpha_{s})\log\hat{p}_{X}(\cdot\mid Y_{\leq t-1},\hat{Y}_{t,...,s-1})$
$-\alpha_{s}\hat{f}_{s+1}(X,[Y_{\leq t-1},\hat{Y}_{t,...,s-1},\cdot]\bigg).$
This expression has no dependence on the first t1t-1 entries $\alpha_{1},...,\alpha_{t-1}$ of $\alpha$, and no dependence on the first t1t-1 entries $\hat{f}_{2},...,\hat{f}_{t}$ of $\hat{f}$, proving the first half of the lemma.
The second half of the lemma follows directly from applying Assumption 5.1, where $\alpha^{\prime},\hat{f}^{\prime}$ and $\alpha,\hat{f}$ can be interchanged by the fact that their future entropies over steps t,,Tt,...,T are the same. ∎
Appendix B Derivations
B.1 Entropy calibration from binary calibration
Recall that for a binary classifier $\hat{f}:\mathcal{X}\to[0,1]$, where $f^{*}:\mathcal{X}\to[0,1]$ denotes the true conditional distribution, binary calibration asks whether the model’s probability corresponds to the actual fraction of ones in reality:
$\mathbb{E}_{X\sim q}\mathbb{E}_{Y\sim f^{*}_{X}}[Y\mid\hat{f}_{X}=p]=p.$
First, note that the right hand side can be replaced by
$\mathbb{E}_{X\sim q}\mathbb{E}_{Y\sim f^{*}_{X}}[Y\mid\hat{f}_{X}=p]=\mathbb{E}_{X\sim q}\mathbb{E}_{\hat{Y}\sim\hat{f}_{X}}[\hat{Y}\mid\hat{f}_{X}=p].$
Next, we can weaken this requirement by making the expectation non-conditional, or
$\mathbb{E}_{X\sim q}\mathbb{E}_{Y\sim f^{*}_{X}}Y=\mathbb{E}_{X\sim q}\mathbb{E}_{\hat{Y}\sim\hat{f}_{X}}\hat{Y}$,
which simply asks whether the overall rate of ones under $\hat{f}$ is the same as the overall rate of ones in reality. The most natural extension of this definition to multiclass calibration is top-class calibration,
$\mathbb{E}_{X\sim q}\mathbb{E}_{Y\sim f^{*}_{X}}\left[\mathbb{1}\left\{Y=\operatorname*{argmax}_{y^{\prime}}\hat{f}_{X}(y^{\prime})\right\}\mid\max_{y^{\prime}}\hat{f}_{X}(y^{\prime})=p\right]=p,$
which states that across all instances where the model assigns $p$ probability to the top class, the actual label should be equal to the top class $p$ fraction of the time on average. Like before, we can replace the right hand side by
$\mathbb{E}_{X\sim q}\mathbb{E}_{Y\sim f^{*}_{X}}\left[\mathbb{1}\left\{Y=\operatorname*{argmax}_{y^{\prime}}\hat{f}_{X}(y^{\prime})\right\}\mid\max_{y^{\prime}}\hat{f}_{X}(y^{\prime})=p\right]$
$\quad=\mathbb{E}_{X\sim q}\mathbb{E}_{\hat{Y}\sim\hat{f}_{X}}\left[\mathbb{1}\left\{\hat{Y}=\operatorname*{argmax}_{y^{\prime}}\hat{f}_{X}(y^{\prime})\right\}\mid\max_{y^{\prime}}\hat{f}_{X}(y^{\prime})=p\right],$
where $Y\sim f^{*}_{X}$ and $\hat{Y}\sim\hat{f}_{X}$ are interchanged. In this expression, the top class probability $\max_{y^{\prime}}\hat{f}_{X}(y^{\prime})$ can be thought of as the confidence of $\hat{f}_{X}$, while the zero-one loss function $\mathbb{1}\left\{\hat{Y}=\operatorname*{argmax}_{y^{\prime}}\hat{f}_{X}(y^{\prime})\right\}$ defines the metric the confidence should be calibrated to — the model’s confidence should correspond to the loss it incurs in reality. For language models, it is natural to replace the zero-one loss with the log loss, which produces the definition
$\mathbb{E}_{X\sim q}\mathbb{E}_{Y\sim f^{*}_{X}}\left[-\log\hat{f}_{X}(Y)\mid H(f_{X})=h\right]$ $=h$
$=\mathbb{E}_{X\sim q}\mathbb{E}_{\hat{Y}\sim\hat{f}_{X}}\left[-\log\hat{f}_{X}(\hat{Y})\mid H(f_{X})=h\right],$
which asks whether the model’s entropy corresponds to the log loss it incurs in reality. We study the unconditional version of this definition
[…]
which simply asks whether the model’s entropy matches its log loss on average. We study unconditional calibration for simplicity, but the same techniques to calibrate unconditionally would likely work for conditional calibration as well if one buckets the inputs $X$ by their entropy $H(f_{X})$.
B.2 Future entropy adjustment from global temperature adjustment
To derive future entropy adjustment from global temperature adjustment, recall that the global temperature adjustment with respect to inverse temperature $\alpha$ (where $\tau=1/(1+\alpha)$) is given by
$p_{\alpha}^{\text{(global)}}(Y_{1},...,Y_{T})=\frac{p(Y_{1},...,Y_{T})^{1+\alpha}}{\sum_{Y^{\prime}\in\mathcal{V}^{T}}p(Y_{1}^{\prime},...,Y_{T}^{\prime})^{1+\alpha}}.$
Factoring this joint distribution into a conditional distribution for each $t$, we have
$\log p_{\alpha}^{\text{(global)}}(Y_{t}\mid Y_{<t})=\log\frac{\sum_{Y_{>t}}p(Y_{<t},Y_{t},Y_{>t})^{1+\alpha}}{\sum_{Y^{\prime}_{t},Y_{>t}}p(Y_{<t},Y^{\prime}_{t},Y_{>t})^{1+\alpha}}.$
Taking the gradient of the log probability with respect to $\alpha$, we have
$\frac{d}{d\alpha}\log p_{\alpha}^{\text{(global)}}(Y_{t}\mid Y_{<t})$ $=\text{softmax}\left\{(1+\alpha)\log p(Y_{<t},Y_{t},Y_{>t}=\cdot)\right\}^{T}\log p(Y_{<t},Y_{t},Y_{>t}=\cdot)$
$-\text{softmax}\left\{(1+\alpha)\log p(Y_{<t},[Y_{t},Y_{>t}]=\cdot)\right\}^{T}\log p(Y_{<t},[Y_{t},Y_{>t}]=\cdot),$
where the first softmax is over $Y_{>t}$ and the second softmax is over both $Y_{t}$ and $Y_{>t}$. Simplifying this expression results in
$=\log p(Y_{t}\mid Y_{<t})+\mathbb{E}_{Y_{>t}\sim p_{\alpha}^{\text{(global)}}(\cdot\mid Y_{\leq t})}\log p(Y_{>t}\mid Y_{\leq t})-\mathbb{E}_{Y_{\geq t}\sim p_{\alpha}^{\text{(global)}}(\cdot\mid Y_{<t})}\log p(Y_{\geq t}\mid Y_{<t}).$
Then, the first-order approximation of $\log p_{\alpha}^{\text{(global)}}(Y_{t}\mid Y_{<t})$ centered around $\alpha=0$ is given by
$\log p_{\alpha}^{\text{(global)}}(Y_{t}\mid Y_{<t})$ $\approx\log p_{\alpha=0}^{\text{(global)}}(Y_{t}\mid Y_{<t})+\alpha\frac{d}{d\alpha}\log p_{\alpha}^{\text{(global)}}(Y_{t}\mid Y_{<t})\Bigg|_{\alpha=0}$
$=\log p(Y_{t}\mid Y_{<t})$
$+\alpha\bigg[\log p(Y_{t}\mid Y_{<t})+\mathbb{E}_{Y_{>t}\sim p_{\alpha=0}^{\text{(global)}}(\cdot\mid Y_{\leq t})}\log p(Y_{>t}\mid Y_{\leq t})$
$-\mathbb{E}_{Y_{\geq t}\sim p_{\alpha=0}^{\text{(global)}}(\cdot\mid Y_{<t})}\log p(Y_{\geq t}\mid Y_{<t})\bigg]$
$=(1+\alpha)\log p(Y_{t}\mid Y_{<t})-\alpha\mathbb{E}_{Y_{>t}\sim p(\cdot\mid Y_{\leq t})}[-\log p(Y_{>t}\mid Y_{\leq t})]+C_{Y_{\leq t}},$
where the final term is constant with respect to $Y_{t}$.
B.3 Future entropy adjustment from MaxEnt RL
The future entropy adjustment can also be derived in the MaxEnt RL framework (ziebart-2008-maximum), where the reward function is given by $r(x,y)=\log\hat{p}_{x}(y)$ with $\hat{p}$ denoting the base model. Specifically, we can write the MaxEnt RL objective as
[…]
Then, the value function for this objective is given by
$V_{X}(Y_{\leq t})$ $=\mathbb{E}_{Y_{>t}\sim\tilde{p}_{X}(Y_{>t}\mid Y_{\leq t})}\log\hat{p}_{X}(Y_{>t}\mid Y_{\leq t}),$
and the Q function is given by
$Q_{X}(Y_{<t},Y_{t})$ $=r_{X}(Y_{t}\mid Y_{<t})+V_{X}(Y_{\leq t})$
$=\log\hat{p}_{X}(Y_{t}\mid Y_{<t})+\mathbb{E}_{Y_{>t}\sim\tilde{p}_{X}(Y_{>t}\mid Y_{\leq t})}\log\hat{p}_{X}(Y_{>t}\mid Y_{\leq t}).$
Using this Q function to define the KL-regularized policy then results in
$\tilde{p}_{\alpha;X}(Y_{t}\mid Y_{<t})$ $\propto\exp\left\{\log\hat{p}_{X}(Y_{t}\mid Y_{<t})+\alpha Q_{X}(Y_{<t},Y_{t})\right\}$
$=\exp\left\{(1+\alpha)\log\hat{p}_{X}(Y_{t}\mid Y_{<t})-\alpha\mathbb{E}_{Y_{>t}\sim\tilde{p}_{\alpha;X}(Y_{>t}\mid Y_{\leq t})}[-\log\hat{p}_{X}(Y_{>t}\mid Y_{\leq t})]\right\},$
which is the future entropy adjustment.
B.4 Scaling in the simplified setting
Recall our simplified setup: the model sees $m$ tokens drawn i.i.d. from an $\alpha$ power law distribution over a vocabulary of size $v$, and it stores the count of each token it sees. At generation time, the model generates a sequence of length $L$ as follows: if the context contains only tokens the model has seen more than once, it behaves normally and produces the next token according to its fitted unigram distribution. But if the context contains at least one token that the model saw only once, then it instead produces the next tokens according to some derailed distribution with entropy larger by some constant $C_{H}$.
First, if the per-step derailing probability $q$ is small, we can compute expected entropy at time $t$ as follows using the binomial approximation:
[…]
$\approx(1-qt)H_{0}+(1-(1-qt))(H_{0}+C_{H})$
$=H_{0}+qtC_{H},$
so the overall miscalibration is given by
$\sum_{t=1}^{L}H_{t}(\hat{p})-H_{0}$ $=\sum_{t=1}^{L}qtC_{H}$
$=qC_{H}\frac{L(L-1)}{2}.$
Next, to characterize the scaling of the expected per-step derailing probability $q$ with respect to dataset size $m$, we first note that
$q=\frac{K_{m,1}}{m},$
where $K_{m,1}$ is a random variable denoting the number of items seen exactly once in the training set of size $m$. Taking the expectation with respect to random draws of the training set, we have
$\mathbb{E}K_{m,1}$ $=\mathbb{E}\sum_{i=1}^{v}\mathbb{1}\{\text{count}_{m}(i)=1\}$
$=\sum_{i=1}^{v}\mathbb{E}\mathbb{1}\{\text{count}_{m}(i)=1\}$
$=\sum_{i=1}^{v}mp_{i}(1-p_{i})^{m-1},$
where $p_{i}=Z/i^{\alpha}$ is the power law probability of the $i$th item, with $Z=\sum_{i=1}^{v}1/i^{\alpha}$ denoting the normalizing constant. Next, taking $v\to\infty$ following the infinite urn setup in good-1953-population; karlin-1967-central, we compute
$\int_{i=1}^{\infty}mp_{i}(1-p_{i})^{m-1}di$ $=\int_{i=1}^{\infty}mZi^{-\alpha}(1-Zi^{-\alpha})^{m-1}di$
$=\frac{1}{\alpha}Z^{\frac{1}{\alpha}}(m-1)^{\frac{1}{\alpha}}\gamma(1-1/\alpha,(m-1)Z),$
where
$\gamma(a,x)$ $=\int_{0}^{x}t^{a-1}e^{-t}dt$
is the lower incomplete gamma function. Taking $m\to\infty$ and using the fact that $\gamma(a,x)\to\Gamma(a)$ for $x\to\infty$, we have that
$\mathbb{E}\frac{K_{m,1}}{m}\sim\frac{1}{\alpha}Z^{\frac{1}{\alpha}}m^{\frac{1}{\alpha}-1}\Gamma(1-1/\alpha),$
as desired. This expression can also be found in Equation 23 of karlin-1967-central.
Appendix C Experimental details
We study four model families (Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B) (qwen2025qwen25technicalreport), Llama 3 (1B, 3B, 8B, 70B) (grattafiori2024llama3herdmodels), Llama 2 (7B, 13B, 70B) (touvron2023llama2openfoundation), and Pythia (410M, 1.4B, 2.8B, 6.9B, 12B) (biderman2023pythiasuiteanalyzinglarge)) applied to the following three datasets:
  1. (a)
    WikiText-103 (merity2016pointersentinelmixturemodels): given 128 tokens of context from a Wikipedia passage, the model is tasked with completing the passage.
  2. (b)
    WritingPrompts (fan-etal-2018-hierarchical): given a writing prompt from the writingprompts subreddit along with 128 tokens of context from a human-written story, the model is tasked with completing the story.
  3. (c)
    CodeContests (li-2022-competition): given a coding problem from one of five websites (e.g. Codeforces) and 128 tokens of context from a human-written solution, the model is tasked with completing the solution.
In each setting, we use 5000 examples and limit the generation to at most 1024 tokens. For generation we use vLLM (kwon2023efficient) with the xFormers attention kernel (xFormers2022) and no quantization, and we use HuggingFace (wolf-etal-2020-transformers) with 4-bit quantization (dettmers2022bnb) to compute logprobs. All experiments are run using PyTorch (pytorch-2019), and all plots are produced using Matplotlib (Hunter:2007). For better readability, the entropy over time plots (Figure 3) are produced with exponential smoothing ($\alpha=0.2$). All experiments are run on 1-4 NVIDIA-A100-SXM4-80GB GPUs, or 1-4 NVIDIA RTX 6000 Ada Generation 49.1GB GPUs.
Appendix D Example generations
In this section, we print excerpts from generations of Qwen2.5-14B applied to WikiText, where we choose three random excerpts each from high, medium, and low entropy buckets (i.e. randomly chosen from the first, 16th, and 32nd entropy buckets). Qualitatively, low entropy generations are either repetitive or contain verbatim copies of the training set, medium entropy generations are high quality, and high entropy generations are incoherent.
Low entropy:
Generation (Entropy=0.548):
- 3.15 Tropical Storm Mischa
- 3.16 Tropical Storm Nigel
- 3.17 Tropical Depression Seventeen
- 3.18 Tropical Storm Patty
- 3.19 Hurricane Rupert
- 3.20 Tropical Storm Sarah
- 3.21 Tropical Storm Tory
- 3.22 Tropical Storm Whitney
- 3.23 Tropical Depression Twenty-two
- 3.24 Tropical Storm Vince
- 3.25 Tropical Storm Wiloma
- 4 Impact
- 5 Season effects
- 6
Generation (Entropy=0.068):
= Performance Review =
"Performance Review" is the eighth episode of the second season of the American comedy television series The Office, and the show’s fourteenth episode overall. It was written by Larry Wilmore and directed by Paul Feig. It first aired on November 15, 2005 on NBC. The episode guest stars Melora Hardin as Jan Levinson.
The series depicts the everyday lives of office employees in the Scranton, Pennsylvania branch of the fictional Dunder Mifflin Paper Company. In this episode, Michael Scott (Steve Carell) conducts job performance reviews with his employees, and struggles
Generation (Entropy=0.817): rebuilt superstructure in a pagoda mast style, displacing her to 32,000 t. These modifications brought her speed down to 21.5 kn (39.8 km/h; 24.7 mph), causing her to be assigned to second-line duties, conducting training operations through 1939.
Following the outbreak of World War II in 1941, Yamashiro took part in the Indochina Incident in late 1940 and early 1941. Shortly before the attack on Pearl Harbor and the Japanese entrance into the war, she conducted
Medium entropy:
Generation (Entropy=2.574): season, any confrontation between contestants or Gleib during a stunt will lead to a girl screaming briefly in anguish before leaving the set for the rest of the game. While only a few teams have reached this period of the game while the game transitioned to a conclusive period, which concluded with Gleib instigating one team to perform another stunt during the bonus time. The winnings range from \$500 to \$5,000 for each round.
Idiotest debuted on August 13, 2014, airing on GSN. An official showcase occurred the following evening on August 19, featuring the
Generation (Entropy=2.557): measured only across the glacier, not along the PIG’s length, and the cross-x data are interpolated.)
In May 2006, scientists found an increase of 1iq Celsius over warm ocean currents surrounding Antarctica – an average of almost .1 iqC warming over the last 100 years.
In 2005, University of Bristol (UK) researchers report, "Recent changes in Antarctic ice streams" and found that "This slowing was likely driven by a piece of ice shelf breaking away from Pine Island Glacier. However, the slow down was only temporary and the effect seemed only to have been temporary."
Generation (Entropy=2.522): premiere of the thirteenth season and then departed permanently, as part of a major overhaul of the cast. She returned in a guest-acting role for the show’s series finale. Abby first appeared on television in June 1979, two years after Jacobs created Dallas, a series about Texas oilmen whose motivations were less virtuous than its male and female leads. He used the same theme for Knots Landing, however, the series was more regulated and politically correct. Whereas Chester’s antagonists were generally viewed as brutish or psychologically ill, Abby was by definition the rich, glamorous and cunning oil tycoon’s daughter;
High entropy:
[…]
Generation (Entropy=5.031): varieties like Lemon Pop, Canadian Grape, Peruvian Peach, Kiwi - tossed with autoimmune syrups and served on the rocks.
16. Maple Syrup recipe
[…]
which are not available anywhere. But just offing Nashville
You need them to find their way.
Polymorphous wonders of the Cross Shades Align by universal rights and bound package for commerciality.
You may be happy Be fit
Appendix E Additional Experiments
Figure 5: MAUVE for excerpts of model generations plotted against the entropy (in nats) of the excerpt, with models colored by size (see Appendix E for the full plots containing all model families). These plots show that sample quality drops when entropy is too high or low.
Figure 6: Entropy calibration error versus log loss for all Llama 2 models: each plot contains per-step-averaged calibration error versus log loss for the base model ($\tau=1.0$) compared to the instruction-tuned version, along with various temperature settings.
Figure 7: Entropy calibration error versus log loss for all Llama 3 models: each plot contains per-step-averaged calibration error versus log loss for the base model ($\tau=1.0$) compared to the instruction-tuned version, along with various temperature settings. Unlike the other model families, instruction tuning on Llama 3 seems to increase calibration error instead of decreasing it. Based on issues that others have also had with these models, we suspect that there might be unresolved issues with the tokenizer configuration. We use the same standard code for all models, and hope to recreate these plots when the issues with the model are resolved.
Figure 8: Entropy calibration error versus log loss for all Qwen2.5 models: each plot contains per-step-averaged calibration error versus log loss for the base model ($\tau=1.0$) compared to the instruction-tuned version, along with various temperature settings.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction are tied to specific experiments and theoretical results in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations are discussed throughout the paper (e.g., when discussing the power law and scaling exponents, the practical feasibility of the algorithm, etc.).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The assumption is stated formally and the appendix contains a full proof.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Please see Appendix C for experimental details.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code will be provided upon acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Please see Appendix C for experimental details.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: Error bars are provided where they would make sense (see, e.g., Figure 5) but omitted when they would not make sense or would clutter the plots visually.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Please see Appendix C.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The paper only uses public datasets and no human subjects. The work is mostly theoretical, and societal implications are discussed below.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: While our work primarily involves analysis and theory, it has implications for downstream tasks like creative writing and code generation. The advancement of language model capabilities in these domains would lead to useful tools, but would also disrupt online communities and people’s livelihoods. We hope that language models can be deployed responsibly, in ways that maintain the health and well-being of the communities they are trained on.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not release any data or models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The paper uses publicly available code packages and datasets and cites them.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 We study the problem of entropy calibration, which asks whether a language model’s entropy over generations matches its log loss on human text.
-Past work found that models are miscalibrated, with entropy per step increasing (and text quality decreasing) as generations grow longer.
-This error accumulation is a fundamental problem in autoregressive models, and the standard solution is to truncate the distribution, which improves text quality at the cost of diversity.
-In this paper, we ask: is miscalibration likely to improve with scale, and is it theoretically possible to calibrate without tradeoffs?
+Past work found that models are miscalibrated, with entropy per step increasing as generations grow longer, due to error accumulation.
+To calibrate the model and improve text quality, it has become standard practice to truncate the distribution, but this approach reduces output diversity, which we would like to avoid.
+Therefore, in this paper, we ask: does miscalibration improve automatically with scale, and if not, is it theoretically possible to calibrate without tradeoffs?
 To build intuition, we first study a simplified theoretical setting to characterize the scaling behavior of miscalibration with respect to dataset size.
-We find that the scaling behavior depends on the power law exponent of the data distribution — in particular, for a power law exponent close to 1, the scaling exponent is close to 0, meaning that miscalibration improves very slowly with scale.
+We find that the rate of scaling depends on the power law exponent of the data distribution — in particular, for a power law exponent close to 1, the scaling exponent is close to 0, meaning that miscalibration improves very slowly with scale.
 Next, we measure miscalibration empirically in language models ranging from 0.5B to 70B parameters.
-We find that the observed scaling behavior is similar to what is predicted by the simplified setting: our fitted scaling exponents for text are close to 0, meaning that larger models accumulate error at a similar rate as smaller ones.
+We find that the observed scaling behavior is similar to what is predicted theoretically: our fitted scaling exponents for text are close to 0, meaning that larger models accumulate error at a similar rate as smaller ones.
 This scaling (or, lack thereof) provides one explanation for why we sample from larger models with similar amounts of truncation as smaller models, even though the larger models are of higher quality.
@@
 1 Introduction
-We study entropy calibration, which asks whether a language model’s entropy over generations matches its log loss on human text.
-This definition is a natural notion of calibration for generative tasks, and is much more challenging than calibration for classification tasks because the output space is exponentially large.
-While we will discuss this definition in more detail later, one basic requirement for calibration is that the model’s entropy per step should be stable over the generation.https://github.com/stevenxcao/entropy-calibration
-Unfortunately, autoregressive language models are not stable.
-Entropy calibration was first studied by braverman2020calibration, who found that language models have entropy per step increasing as generations grow longer.
-This entropy blowup is accompanied by an increase in generation errors, as is also observed in basu2021mirostatneuraltextdecoding.
-While entropy calibration specifically is not well studied, generation instability more broadly has been the subject of many papers (williams1989learning; Ranzato2015SequenceLT; Welleck2019NeuralTG).
-From this line of work has emerged a suite of distribution truncation techniques, which have become standard practice in language model sampling (fan-etal-2018-hierarchical; holtzman2020curious; hewitt-etal-2022-truncation).
-These techniques suppress low probability tokens to improve quality at the cost of diversity (hashimoto-etal-2019-unifying; zhang-etal-2021-trading).
-It is not fully satisfying that to make generation stable, we must sacrifice diversity.
-Diversity is especially important for difficult tasks where we must aggregate multiple answers (wang2024mixtureofagentsenhanceslargelanguage; brown2024largelanguagemonkeysscaling), as well as for synthetic data generation, which has seen a resurgence of interest as the community has begun worrying about running out of internet data (wang-etal-2023-self-instruct; gunasekar2023textbooksneed; maini-etal-2024-rephrasing).
-Therefore, it is natural to ask: do we expect generation stability to improve with scale?
+We study entropy calibration, which asks whether a language model’s entropy over generations matches its log loss on human text.https://github.com/stevenxcao/entropy-calibration This definition is a natural notion of calibration for generative tasks, and is more challenging than calibration for classification tasks because the output space is exponentially large.
+Braverman et al.
+(2020) were the first to study entropy calibration in language models, and they found that models are miscalibrated: entropy per step for model generations increases with the length of the document, in contrast with log loss on human text which is roughly flat over the length of the document (Genzel and Charniak, 2002; Verma et al., 2023).
+Entropy calibration can be thought of as a quantitative formalization of the well-known “error accumulation” or “teacher forcing” problem: entropy rises when the model generates erroneous tokens which are fed back into the context, derailing the generation (Williams and Zipser, 1989; Ranzato et al., 2016; Basu et al., 2021; Hewitt et al., 2022, also see Appendix E).
+Therefore, we study entropy calibration to gain fundamental insights into improving generation quality.
+To correct error accumulation and calibrate the model, it has become standard practice to truncate the next token distribution (Fan et al., 2018; Holtzman et al., 2020; Hewitt et al., 2022), suppressing low probability tokens to improve quality at the cost of diversity (Hashimoto et al., 2019; Zhang et al., 2021).
+This solution is not satisfying: diversity is especially important for difficult tasks where we must aggregate multiple answers (Wang et al., 2024; Brown et al., 2024), as well as for synthetic data generation, which has seen a resurgence of interest as the community has begun worrying about running out of internet data (Wang et al., 2023; Gunasekar et al., 2023; Maini et al., 2024).
+Therefore, it is natural to ask: do we expect miscalibration to improve with scale?
 If not, is it at least theoretically possible to calibrate without sacrificing diversity?
@@
 This unfamiliar token then derails subsequent steps when it is fed back into the context autoregressively.
-Drawing on classic results, we calculate a scaling exponent capturing how quickly the probability of generating a rare token decreases with the number of training examples (good-1953-population; karlin-1967-central).
-We find that this exponent depends on how heavy-tailed the data distribution is: in particular, for power law exponents close to 1, as is typical for human text (zipf-1936-psycho; zipf-1949-human), the scaling exponent is close to 0.
+Drawing on classic results, we calculate a scaling exponent capturing how quickly the probability of generating a rare token decreases with the number of training examples (Good, 1953; Karlin, 1967).
+We find that this exponent depends on how heavy-tailed the data distribution is: in particular, for power law exponents close to 1, as is typical for human text (Zipf, 1936, 1949), the scaling exponent is close to 0.
 Therefore, this setup predicts that stability in generation improves very slowly with scale.
@@
 We find that much like distribution truncation, instruction tuning reduces entropy at the cost of increased log loss, with the largest models now having entropy too low.
-This tradeoff relates to past work which found that alignment degrades model capabilities, a phenomenon known as the alignment tax (ouyang2022training; bai2022helpful; lin-etal-2024-mitigating).
+This tradeoff relates to past work which found that alignment degrades model capabilities, a phenomenon known as the alignment tax (Ouyang et al., 2022; Bai et al., 2022; Lin et al., 2024).
 Given that all known mitigations increase the model’s log loss, is it even possible in theory to calibrate without this tradeoff?
@@
 2 Preliminaries
-We first review key definitions and properties for entropy calibration, introduced in braverman2020calibration.
+We first review key definitions and properties for entropy calibration, introduced in Braverman et al.
+(2020).
 Our setup is as follows: we are given prompts $X\in\mathcal{X}$ drawn from some prompt distribution $X\sim q$, and responses $Y\in\mathcal{Y}$ drawn from the true conditional distribution $Y\sim p^{*}_{X}$.
@@
 This definition mirrors that of binary calibration, and we derive this connection more formally in Appendix B.
-Qualitatively, if a model is underconfident, then its generations have too much entropy and appear incoherent; if it is overconfident, then its generations have too little entropy and appear repetitive (braverman2020calibration; basu2021mirostatneuraltextdecoding); see Appendix E for a replication of this finding and Appendix D for examples.
+Qualitatively, if a model is underconfident, then its generations have too much entropy and appear incoherent; if it is overconfident, then its generations have too little entropy and appear repetitive (Braverman et al., 2020; Basu et al., 2021); see Appendix E for a replication of this finding and Appendix D for examples.
 Entropy calibration is then the problem of adjusting the entropy to be just right.
-Empirically, braverman2020calibration found that neural autoregressive language models have entropy too high: entropy per step matches the log loss at earlier steps but increases as the generation grows.
+Empirically, Braverman et al.
+(2020) found that neural autoregressive language models have entropy too high: entropy per step matches the log loss at earlier steps but increases as the generation grows.
 Why does entropy per step grow with the length of the generation?
@@
 In contrast, models must generate multiple tokens at deployment time, which they do by producing one token at a time and taking their own production as subsequent input.
-Therefore, even models with very low single-step error can degrade over multiple steps as they take their own slightly erroneous outputs as input and accumulate errors (see, e.g., Ranzato2015SequenceLT, Welleck2019NeuralTG, holtzman2020curious for error accumulation in language modeling; and daume2009search, ross2010efficient, ross2011reduction for imitation learning).
-This intuition is formalized in the context of entropy calibration in Corollary 4.2 of braverman2020calibration, which states that for a model with $\varepsilon$ KL divergence to the true distribution, the entropy at step $t$ can deviate as much as $\varepsilon+\sqrt{\varepsilon t}$ from the log loss, growing with $t$.
+Therefore, even models with very low single-step error can degrade over multiple steps as they take their own slightly erroneous outputs as input and accumulate errors (see, e.g., Ranzato et al.
+(2016), Welleck et al.
+(2020), Holtzman et al.
+(2020) for error accumulation in language modeling; and Daumé et al.
+(2009), Ross and Bagnell (2010), Ross et al.
+(2011) for imitation learning).
+This intuition is formalized in the context of entropy calibration in Corollary 4.2 of Braverman et al.
+(2020), which states that for a model with $\varepsilon$ KL divergence to the true distribution, the entropy at step $t$ can deviate as much as $\varepsilon+\sqrt{\varepsilon t}$ from the log loss, growing with $t$.
 How does one calibrate the entropy?
@@
 In practice, practitioners use a number of distribution truncation methods, each of which uses a different heuristic to suppress low probability tokens in each generation step.
-Some standard methods include temperature reduction, top-k sampling (fan-etal-2018-hierarchical), top-p sampling  (holtzman2020curious), and min-p sampling (hewitt-etal-2022-truncation).
-These methods improve text quality at the cost of diversity (hashimoto-etal-2019-unifying; zhang-etal-2021-trading; pillutla2021mauve; welleck2024decodingmetagenerationinferencetimealgorithms).
-Following hashimoto-etal-2019-unifying, we define a model’s diversity to be its log loss on reference documents.
+Some standard methods include temperature reduction, top-k sampling (Fan et al., 2018), top-p sampling  (Holtzman et al., 2020), and min-p sampling (Hewitt et al., 2022).
+These methods improve text quality at the cost of diversity (Hashimoto et al., 2019; Zhang et al., 2021; Pillutla et al., 2021; Welleck et al., 2024).
+Following Hashimoto et al.
+(2019), we define a model’s diversity to be its log loss on reference documents.
 The intuition behind this definition is that log loss (also known as cross entropy or forward KL) is a coverage metric: to attain low log loss, the model must “cover” as much as the reference distribution as possible.
-Our goal, then, is to calibrate entropy to match log loss without also causing the log loss to also increase.
-In theory, braverman2020calibration show that one can calibrate entropy while preserving log loss via globally normalized temperature scaling, where the adjusted model is given by $\hat{p}_{\tau}(y_{1},...,y_{L})\propto\hat{p}(y_{1},...,y_{L})^{1/\tau}$.
+Our goal, then, is to calibrate entropy to match log loss without also causing the log loss to increase.
+In theory, Braverman et al.
+(2020) show that one can calibrate entropy while preserving log loss via globally normalized temperature scaling, where the adjusted model is given by $\hat{p}_{\tau}(y_{1},...,y_{L})\propto\hat{p}(y_{1},...,y_{L})^{1/\tau}$.
 Unfortunately, this adjustment is intractable to compute because it involves normalizing over the entire output space.
@@
 To gain intuition, we first explore this question in a simplified theoretical setting.
-We define the setup to capture the following hypothesis regarding error accumulation (see, e.g., hewitt-etal-2022-truncation): because the language distribution is heavy-tailed, the model must assign non-zero probability to a large number of rare tokens when fitting the data.
+We define the setup to capture the following hypothesis regarding error accumulation (see, e.g., Hewitt et al.
+(2022)): because the language distribution is heavy-tailed, the model must assign non-zero probability to a large number of rare tokens when fitting the data to achieve low log loss.
 However, if it happens to generate one such rare token, the model derails when that token is fed back into the context autoregressively, leading to a jump in entropy.
@@
 In this setting, the expected entropy per step grows with slope proportional to the probability of emitting a rare token (see Appendix B).
-Computing the rare token mass in power law data is a classic problem, and we can compute the asymptotic scaling exponent with respect to the number of training examples $m$ as follows (good-1953-population; karlin-1967-central):
+Computing the rare token mass in power law data is a classic problem, and we can compute the asymptotic scaling exponent with respect to the number of training examples $m$ as follows (Good, 1953; Karlin, 1967):
 Proposition 3.1 (informal).
@@
 We provide the derivation in the appendix.
-The key takeaway from this proposition is that the derailing probability scales as $m^{1/\alpha-1}$, which is very slow if the power law exponent $\alpha$ is close to $1$, as is typical for text (zipf-1936-psycho; zipf-1949-human).
+The key takeaway from this proposition is that the derailing probability scales as $m^{1/\alpha-1}$, which is very slow if the power law exponent $\alpha$ is close to $1$, as is typical for text (Zipf, 1936, 1949).
 The reason for this slow scaling is that as $m$ increases, there are always more rare items to be sampled from the tail of the distribution.
@@
 These phenomena are then memorized, and potentially generated at deployment, derailing the model.
-While asymptotic analysis gives us a clean expression, we can also estimate the scaling exponent in simulation for finite values of $m$ and $v$ (see Figure 1).
-We find that the actual slopes are close to the asymptotic expression, up to $m\approx v/3$.
+While asymptotic analysis gives us a clean expression, we can also estimate the scaling exponent in simulation for finite values of $m$ and $v$.
+We find that the non-asymptotic simulated slopes are close to the asymptotic expression as long is $m$ is smaller than $v/3$ (see Appendix E).
 We also calculate the power law exponent for our three datasets, finding that it is around 1 for WikiText and WritingPrompts and $1.5$ for CodeContests, which predicts slow scaling for the first two datasets and slow-to-moderate scaling for the third.
-Figure 1: Left: the expected total mass of tokens seen exactly once, given m samples from a power law distribution over a vocabulary of size $v$, for three settings of the power law exponent $\alpha=1.0,1.25,1.5$.
-Their relationship is roughly log-log linear up to $m\approx v/3$, with slope slightly steeper than the asymptotic expression of $1/\alpha-1$.
-Right: log frequency versus log rank of the top 5000 unigrams in three datasets.
-The power law exponent $\alpha$, given by the slope of each curve, is close to 11 for WikiText and WritingPrompts, while it is $1.5$ for CodeContests, suggesting that text has heavier tails than code.
-Together, these plots suggest that the singleton mass should decay more slowly with $m$ for WikiText and WritingPrompts than for CodeContests.
+Figure 1: Log calibration error versus log model size for four model families and three datasets.
+We find that the scaling laws fit relatively well, suggesting that the relationship between calibration and scale is predictable.
+Furthermore, while there is variation between model families, the scaling exponents for each dataset are somewhat close to those predicted by theory (WikiText: $0.089$, WritingPrompts: 0.10-0.10, CodeContests: 0.33-0.33), with heavier-tailed datasets having slower scaling.
 4 Experiments: Miscalibration in Large Language Models
 Next, we measure miscalibration empirically in large language models.
-We study four model families (Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B) (qwen2025qwen25technicalreport), Llama 3 (1B, 3B, 8B, 70B) (grattafiori2024llama3herdmodels), Llama 2 (7B, 13B, 70B) (touvron2023llama2openfoundation), and Pythia (410M, 1.4B, 2.8B, 6.9B, 12B) (biderman2023pythiasuiteanalyzinglarge)) applied to the three datasets listed below.
+We study four model families (Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B) (Qwen et al., 2025), Llama 3 (1B, 3B, 8B, 70B) (Grattafiori et al., 2024), Llama 2 (7B, 13B, 70B) (Touvron et al., 2023), and Pythia (410M, 1.4B, 2.8B, 6.9B, 12B) (Biderman et al., 2023)) applied to the three datasets listed below.
 In each setting, we use 5000 examples and limit samples to 1024 tokens; see Appendix C for more experimental details.
@@
   1. (a)
-    WikiText-103 (merity2016pointersentinelmixturemodels): given 128 tokens of context from a Wikipedia passage, the model is tasked with completing the passage.
+    WikiText-103 (Merity et al., 2017): given 128 tokens of context from a Wikipedia passage, the model is tasked with completing the passage.
   2. (b)
-    WritingPrompts (fan-etal-2018-hierarchical): given a prompt from r/writingprompts along with 128 tokens of context from a human-written story, the model is tasked with completing the story.
+    WritingPrompts (Fan et al., 2018): given a prompt from r/writingprompts along with 128 tokens of context from a human-written story, the model is tasked with completing the story.
   3. (c)
-    CodeContests (li-2022-competition): given a coding problem from one of five websites and 128 tokens of context from a human-written solution, the model is tasked with completing the solution.
+    CodeContests (Li et al., 2022): given a coding problem from one of five websites and 128 tokens of context from a human-written solution, the model is tasked with completing the solution.
 4.1 Miscalibration scaling in base models
-Figure 2: Log calibration error versus log model size for four model families and three datasets.
-We find that the scaling laws fit relatively well, suggesting that the relationship between calibration and scale is predictable.
-Furthermore, while there is variation between model families, the scaling exponents for each dataset are somewhat close to those predicted by theory (WikiText: $0.089$, WritingPrompts: 0.10-0.10, CodeContests: 0.33-0.33), with heavier-tailed datasets having slower scaling.
-Past work has found that many model capabilities improve predictably with model size, with task loss and model size following a linear relationship when plotted on a log scale (kaplan2020scalinglawsneurallanguage; hoffmann2022trainingcomputeoptimallargelanguage).
+Figure 2: Entropy for each generation step (solid) and log loss for each token in the ground truth (dashed), for each dataset (columns) and each model family (rows), with models colored by size.
+Models have entropy much higher than their log loss, with the gap growing with the number of generation steps, a sign of error accumulation.
+For the text datasets, models of different sizes seem to be similarly miscalibrated, while for code the degree of miscalibration seems to improve with size.
+Past work has found that many model capabilities improve predictably with model size, with task loss and model size following a linear relationship when plotted on a log scale (Kaplan et al., 2020; Hoffmann et al., 2022).
 We use a similar methodology to study the relationship between entropy miscalibration and model size.
@@
 Overall, these plots suggest that miscalibration in text generation improves very slowly with scale: a scaling exponent of 0.10-0.10 means that to reduce calibration error by a factor of $10$, dataset size must increase by a factor of $10^{10}$.
-Figure 3: Entropy for each generation step (solid) and log loss for each token in the ground truth (dashed), for each dataset (columns) and each model family (rows), with models colored by size.
-Models have entropy much higher than their log loss, with the gap growing with the number of generation steps, a sign of error accumulation.
-For the text datasets, models of different sizes seem to be similarly miscalibrated, while for code the degree of miscalibration seems to improve with size.
 4.2 Entropy over time
+Figure 3: Entropy calibration error versus log loss for base Qwen2.5 (1.5B, 7B, 72B) compared to the instruction-tuned versions, along with various temperature settings (please see Appendix E for all model sizes).
+Positive calibration error means that the model’s entropy is higher than its log loss, while negative means that its entropy is lower than its log loss.
+We find that each modification of the base model reduces entropy while increasing log loss, calibrating at the cost of diversity.
 To gain a more fine-grained understanding of entropy blowup, we also produce entropy over time plots for each model and each dataset, as shown in Figure 2.
@@
 Past papers use a model’s log loss to estimate the actual entropy of the underlying text, as the former is an upper bound for the latter that grows tighter if the model is more accurate.
-This part of the plot replicates past findings that the entropy per step of human text is stable over time, also known as the entropy rate constancy principle (genzel-charniak-2002-entropy; verma-etal-2023-revisiting).
+This part of the plot replicates past findings that the entropy per step of human text is stable over time, also known as the entropy rate constancy principle (Genzel and Charniak, 2002; Verma et al., 2023).
 On the other hand, unlike human text, the entropy per step of model generations is not stable and instead increases over time.
@@
 For CodeContests, the slopes decrease with model size, visually confirming that there is slow-to-moderate scaling.
-Figure 4: Entropy calibration error versus log loss for base Qwen2.5 (1.5B, 7B, 72B) compared to the instruction-tuned versions, along with various temperature settings (please see Appendix E for all model sizes).
-Positive calibration error means that the model’s entropy is higher than its log loss, while negative means that its entropy is lower than its log loss.
-We find that each modification of the base model reduces entropy while increasing log loss, calibrating at the cost of diversity.
 4.3 Calibration-diversity tradeoffs
@@
 We then plot entropy calibration error against log loss, where each setting of the model is one point on the plot, as shown in Figure 3.
-First, we find that reducing temperature below $1$ reduces entropy but increases log loss, replicating similar findings in past work (hashimoto-etal-2019-unifying).
+First, we find that reducing temperature below $1$ reduces entropy but increases log loss, replicating similar findings in past work (Hashimoto et al., 2019).
 Furthermore, the temperature attaining zero calibration error is similar across model sizes, which makes sense given that they are similarly miscalibrated.
-We find that instruction tuning also reduces entropy while increasing log loss, which is consistent with past work showing that instruction tuning harms diversity (ghosh2024closer).
+We find that instruction tuning also reduces entropy while increasing log loss, which is consistent with past work showing that instruction tuning harms diversity (Ghosh et al., 2024).
 Unlike temperature scaling, the magnitude of the effect varies across model sizes, with larger models experiencing both a larger reduction in entropy and larger increase in log loss.
@@
 Why does future entropy preserve log loss?
-Future entropy adjustment can be derived as a first-order approximation of globally normalized temperature adjustment; we provide this derivation in Appendix B, along with a derivation in the MaxEnt RL framework (ziebart-2008-maximum).
-Global temperature adjustment attains calibration as long as the gradient of the log loss with respect to temperature is small (braverman2020calibration), which is a first-order condition.
+Future entropy adjustment can be derived as a first-order approximation of globally normalized temperature adjustment; we provide this derivation in Appendix B, along with a derivation in the MaxEnt RL framework (Ziebart et al., 2008).
+Global temperature adjustment attains calibration as long as the gradient of the log loss with respect to temperature is small (Braverman et al., 2020), which is a first-order condition.
 Then, intuitively, a first-order approximation of global temperature scaling should preserve this property.
@@
 One other point to note is that our analysis holds for any approximation of the future entropy that attains error $\delta$, with worse approximations just weakening the calibration error guarantee.
-For example, one could use the one-step future entropy (braverman2020calibration), or truncate to $k$ steps instead.
+For example, one could use the one-step future entropy (Braverman et al., 2020), or truncate to $k$ steps instead.
 We hope that our theory, which establishes future entropy as the target to approximate, guides future work to achieve better quality-diversity tradeoffs than existing approaches.
 6 Additional Related Work
-Calibration is most commonly studied in binary and multiclass classification, with some classic algorithms including binning, Platt scaling, and isotonic regression (PlattProbabilisticOutputs1999; zadrozny2002transforming; guo2017on; kumar2019verified).
-In the language modeling setting, liang2023holisticevaluationlanguagemodels evaluate the calibration of language models prompted to perform a wide range of classification tasks, finding that models are almost always miscalibrated and overconfident.
+Calibration is most commonly studied in binary and multiclass classification, with some classic algorithms including binning, Platt scaling, and isotonic regression (Platt, 1999; Zadrozny and Elkan, 2002; Guo et al., 2017; Kumar et al., 2019).
+In the language modeling setting, Liang et al.
+(2023) evaluate the calibration of language models prompted to perform a wide range of classification tasks, finding that models are almost always miscalibrated and overconfident.
 In such a setting, one can simply apply standard calibration techniques to adjust the model’s outputted probabilities.
 More challenging is linguistic calibration, where models appear overconfident in the language they use to answer a question.
-To address this problem, past works propose techniques based on controllable generation and reinforcement learning (mielke-etal-2022-reducing; band2024linguistic).
-Finally, the term “calibration” is also used to describe the procedure of eliminating the model’s innate bias toward certain tokens when doing in-context learning, to improve task performance (pmlr-v139-zhao21c).
+To address this problem, past works propose techniques based on controllable generation and reinforcement learning (Mielke et al., 2022; Band et al., 2024).
+Finally, the term “calibration” is also used to describe the procedure of eliminating the model’s innate bias toward certain tokens when doing in-context learning, to improve task performance (Zhao et al., 2021).
 All of these settings are distinct from our setting, which studies the calibration of a model’s entropy over an entire generation, and whose related work we discuss in Section 2.
@@
 B.3 Future entropy adjustment from MaxEnt RL
-The future entropy adjustment can also be derived in the MaxEnt RL framework (ziebart-2008-maximum), where the reward function is given by $r(x,y)=\log\hat{p}_{x}(y)$ with $\hat{p}$ denoting the base model.
+The future entropy adjustment can also be derived in the MaxEnt RL framework (Ziebart et al., 2008), where the reward function is given by $r(x,y)=\log\hat{p}_{x}(y)$ with $\hat{p}$ denoting the base model.
 Specifically, we can write the MaxEnt RL objective as
@@
 where $p_{i}=Z/i^{\alpha}$ is the power law probability of the $i$th item, with $Z=\sum_{i=1}^{v}1/i^{\alpha}$ denoting the normalizing constant.
-Next, taking $v\to\infty$ following the infinite urn setup in good-1953-population; karlin-1967-central, we compute
+Next, taking $v\to\infty$ following the infinite urn setup in Good (1953); Karlin (1967), we compute
 $\int_{i=1}^{\infty}mp_{i}(1-p_{i})^{m-1}di$ $=\int_{i=1}^{\infty}mZi^{-\alpha}(1-Zi^{-\alpha})^{m-1}di$
@@
 $\mathbb{E}\frac{K_{m,1}}{m}\sim\frac{1}{\alpha}Z^{\frac{1}{\alpha}}m^{\frac{1}{\alpha}-1}\Gamma(1-1/\alpha),$
-as desired. This expression can also be found in Equation 23 of karlin-1967-central.
+as desired. This expression can also be found in Equation 23 of Karlin (1967).
 Appendix C Experimental details
-We study four model families (Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B) (qwen2025qwen25technicalreport), Llama 3 (1B, 3B, 8B, 70B) (grattafiori2024llama3herdmodels), Llama 2 (7B, 13B, 70B) (touvron2023llama2openfoundation), and Pythia (410M, 1.4B, 2.8B, 6.9B, 12B) (biderman2023pythiasuiteanalyzinglarge)) applied to the following three datasets:
+We study four model families (Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B) (Qwen et al., 2025), Llama 3 (1B, 3B, 8B, 70B) (Grattafiori et al., 2024), Llama 2 (7B, 13B, 70B) (Touvron et al., 2023), and Pythia (410M, 1.4B, 2.8B, 6.9B, 12B) (Biderman et al., 2023)) applied to the following three datasets:
   1. (a)
-    WikiText-103 (merity2016pointersentinelmixturemodels): given 128 tokens of context from a Wikipedia passage, the model is tasked with completing the passage.
+    WikiText-103 (Merity et al., 2017): given 128 tokens of context from a Wikipedia passage, the model is tasked with completing the passage.
   2. (b)
-    WritingPrompts (fan-etal-2018-hierarchical): given a writing prompt from the writingprompts subreddit along with 128 tokens of context from a human-written story, the model is tasked with completing the story.
+    WritingPrompts (Fan et al., 2018): given a writing prompt from the writingprompts subreddit along with 128 tokens of context from a human-written story, the model is tasked with completing the story.
   3. (c)
-    CodeContests (li-2022-competition): given a coding problem from one of five websites (e.g.
+    CodeContests (Li et al., 2022): given a coding problem from one of five websites (e.g.
 Codeforces) and 128 tokens of context from a human-written solution, the model is tasked with completing the solution.
 In each setting, we use 5000 examples and limit the generation to at most 1024 tokens.
-For generation we use vLLM (kwon2023efficient) with the xFormers attention kernel (xFormers2022) and no quantization, and we use HuggingFace (wolf-etal-2020-transformers) with 4-bit quantization (dettmers2022bnb) to compute logprobs.
-All experiments are run using PyTorch (pytorch-2019), and all plots are produced using Matplotlib (Hunter:2007).
+For generation we use vLLM (Kwon et al., 2023) with the xFormers attention kernel (Lefaudeux et al., 2022) and no quantization, and we use HuggingFace (Wolf et al., 2020) with 4-bit quantization (Dettmers et al., 2022) to compute logprobs.
+All experiments are run using PyTorch (Paszke et al., 2019), and all plots are produced using Matplotlib (Hunter, 2007).
 For better readability, the entropy over time plots (Figure 2) are produced with exponential smoothing ($\alpha=0.2$).
@@
 Appendix E Additional Experiments
+Figure 4: Left: the expected total mass of tokens seen exactly once, given m samples from a power law distribution over a vocabulary of size $v$, for three settings of the power law exponent $\alpha=1.0,1.25,1.5$.
+Their relationship is roughly log-log linear up to $m\approx v/3$, with slope slightly steeper than the asymptotic expression of $1/\alpha-1$.
+Right: log frequency versus log rank of the top 5000 unigrams in three datasets.
+The power law exponent $\alpha$, given by the slope of each curve, is close to 11 for WikiText and WritingPrompts, while it is $1.5$ for CodeContests, suggesting that text has heavier tails than code.
+Together, these plots suggest that the singleton mass should decay more slowly with $m$ for WikiText and WritingPrompts than for CodeContests.
 Figure 5: MAUVE for excerpts of model generations plotted against the entropy (in nats) of the excerpt, with models colored by size (see Appendix E for the full plots containing all model families).